### 06 Mar 2024 v2

# The Still Secret Ballot:

Abstract
After an election, should officials release an electronic record of each ballot? The release of ballots could bolster the legitimacy of the result. But it may also facilitate vote revelation, where an analyst unravels the secret ballot by uniquely linking votes on an anonymous ballot to the voter’s name and address in the public voter file. We first provide a theoretical model of how vote revelation could occur under various election-reporting regimes. Perhaps counterintuitively, releasing ballot records is no more revelatory than the typical practice of releasing aggregate vote tallies. We then present the first empirical evaluation of vote revelation, using the 2020 election in Maricopa County, Arizona, as a case study. For 99.8% of voters, the release of ballot records led to no revelation of any vote choice. We conclude the ballot can be both public and still as secret as it is under the current reporting practices.
Introduction
Democracies face a difficult tradeoff in election reporting. After holding an election, election officials must report results at a level of granularity sufficient to bolster the legitimacy of the vote count while avoiding revealing individual vote choices. Where to strike the balance between transparency and privacy is an increasingly pressing question: A growing swath of voters who do not believe their votes are counted accurately are seeking “citizen audits” of elections [61, 56, 10].
States and localities vary in how they report election results. North Carolina offers a particularly illuminating example of the transparency-privacy tradeoff. The State Board of Elections is obligated to both report granular election results by precinct and vote method and protect the state constitutional right to a secret ballot (N.C. Gen. Stat. §163-132.5G). To strike that balance, the board adds or subtracts an undisclosed random number to some reported election results. The board is concerned with what we term vote revelation: an analyst uniquely linking the anonymous vote choices reflected in election results with the public record of voters’ name and addresses. However, as the board acknowledges, the random noise means that a candidate who appears to have won the most votes in the published noised election results may not be the one who was the actual certified winner. [^1] The state website explains that noising may at times have the appearance “of changing the outcome of very close contests.” See https://perma.cc/URR8-Z5YQ and https://perma.cc/F646-EG4Y.
The downside of adding noise to election results is that it makes the results less accurate and possibly misleading [37]. Such privacy-protecting manipulations can also impede the application of election forensics to detect genuine electoral manipulation or malfunction [[, see, e.g.,]]mebane2008election, cantu2014identifying, callen2015institutional, cottrell2018exploration. In short, granular election results are important in order for elections to be accepted as legitimate. At the same time, voter privacy contributes to the legitimacy of the result too. In fact, the secret ballot was adopted to reduce corruption and voter intimidation [43, 48].
This study presents the first empirical estimates of privacy loss under different election reporting regimes, including the public release of anonymous ballots. We use the complete set of individual-level ballot records, known as cast vote records, in Maricopa County, Arizona, as the primary case study. The dataset allows us to compute the extent of potential privacy violations under various reporting regimes, holding voting behavior constant. We estimate that the public release of cast vote records could reveal about 0.0003% of votes for President in the November 2020 election, among the over 2 million voters we consider. Moreover, for more than 99.8% of voters, no revelation of any vote choice in any contest was possible across the 60-70 contests that they voted on.
We conclude that the ballot can still be considered secret even if cast vote records are subsequently made public, for two reasons we discuss in detail. First, the fraction of potential votes revealed is small enough that privacy concerns can be addressed without withholding all cast vote records. Second, the release of cast vote records is no more revelatory than the current practice of reporting aggregate vote tallies by precinct and vote method. The ballot is thus still secret in the sense that the release of cast vote records does not substantially change the privacy risks that already exist when jurisdictions release granular aggregate election results.
Our argument thus differs from the privacy-centric view put forth in academia and the courts. For example, [2] argue against releasing images of each ballot cast because of possible risks to the secret ballot. Further, the top state court in New York ruled that anonymous individual ballot records were not public record in part because they might reveal vote choices. The Governor of Arizona recently vetoed a bipartisan bill to release cast vote records for a similar reason [36]. But neither [2] nor the New York court nor the governor of Arizona estimated the extent of vote revelation.
Our study offers three contributions. First, we propose a way for election administration to incorporate concepts from the burgeoning field of privacy. In the last two decades, a movement largely led by computer scientists has shown how ostensibly private datasets can have their privacy undone by data triangulation [1, 62, 23, 68]. But importing the dominant approach in data privacy to the electoral context is not straightforward. Election data has several distinct features compared to the commercial or medical datasets commonly used in privacy studies: ballots themselves do not contain personally identifying information, the certified result must provide finality, and votes should be anonymous not just to the public, but also to election officials.
Second, using our framework, we provide the first empirical assessment of the privacy costs of transparent election results. Although a limited set of academic studies and assessments by expert commissions has already pointed to the potential for vote revelation from releasing cast vote records, those studies do not go on to measure the extent of the problem in an actual election [2, 50, 8], or, if they do, only measure a special case of vote revelation [13]. Our study clarifies that the reporting of election results should only lead to what we term vote revelation when there is a unanimous result in a reporting unit (for example, in a precinct of 10 voters, all 10 of them vote for the Republican candidate). Further, it shows how the release of cast vote records is no more revelatory than the release of election results aggregated by the partially identifying information (what we call quasi-identifiers) available in cast vote records.
Third, and finally, the extent of revelation is directly relevant to current litigation about whether cast vote records should be public record for transparency or kept confidential because of privacy. The extent of revelation is also pertinent to policy decisions about how to best strike the difficult balance between transparency and privacy in election reporting. We estimate that less than 3,500 out of over 2 million Maricopa County voters could have any of their votes revealed. Further, when votes are revealed, they are disproportionately provisionally cast ballots, which only comprise 0.1 percent of total ballots cast, and are usually votes for candidates who are popular in the area in which the voter lives. These findings help inform whether a particular response—ranging from not releasing any individual ballots to aggregating, suppressing, or adding random noise to results—is proportionate to the potential privacy loss.
Current Challenges in Election Reporting
Across the country, election administrators and state officials alike are currently grappling with how to best report election results. In general, election administrators have typically tallied individual ballots and reported aggregate results. For example, almost every state reported results for the 2016 and 2020 presidential elections results by precinct [6]. But vote tabulation technology has been changing. Most modern voting systems can also preserve anonymous, individual ballot records in addition to tallying aggregate results. The changing technology has made clear the tradeoff between voter privacy and public transparency in election reporting.
The rise of election skepticism has triggered increased interest in individual ballot records. A number of election officials now affirmatively release these ballot records, including the states of Alaska and Maryland, as well as the counties of San Francisco, California, Dane County, Wisconsin, and Leon County, Florida [4]. But many other election officials have prohibited the release of ballot records. For example, a review of recent news reports reveals that election officials in Indiana, Missouri, and South Carolina have all denied requests for ballot records [58, 70, 38].
Election officials have come to different conclusions in part because of an empirical disagreement about the consequences of public disclosure [24]. For example, the Alaska Division of Elections has released individual ballots because “ballots are anonymous and not connected to voters” [7]. Similarly, Dane County has noted that “there [is] nothing linking the data to individual voters.” In contrast, the South Carolina Election Commission has denied public records requests because “the release of [ballot records] would likely lead to [the] identification of voters” [39]. Similarly, the Indiana Secretary of State Election Division determined that disclosure “would identify each voter” [54]. The conflict has come to bear within states too. For example, Arizona’s Democratic Secretary of State recently supported a bill to release individual ballots because it “would help restore trust in [the] election process” [26], but the Democratic governor ultimately vetoed it, arguing it “threatens anonymity” [36].
State courts have also reached conflicting conclusions about the release of ballot records because of the same empirical disagreement. Their legal analysis focuses on the specifics of state public records law. In general, under the typical state public records law, the release of ballots depends on whether public disclosure would violate the state right to a secret ballot. For example, a Pennsylvania court recently held that the release of ballot records does not violate the secret ballot provision in the state’s constitution because disclosure will not reveal voters’ selections (Honey v. Lycoming County, No. 22-cv-115 (Pa. Commw. Ct. Dec. 15, 2022)). In contrast, the New York Court of Appeals, the highest court in the state, has come to the opposite conclusion. (Kosmider v. Whitney, 132 N.E.3d 592 (N.Y. 2019)). There is also active litigation in other states. For example, a Missouri county clerk recently filed suit to seek clarification about the public status of individual ballot records (Petition for Declaratory Judgement, Schoeller v. Huddleston, No. 2231-CC00975 (Greene County, Mo. Cir. Ct. 2022)).
Vote Revelation
The lack of consensus about the release of individual ballot records reflects, at least in part, the lack of empirical research. Election officials and courts alike have been left to navigate the new technological landscape with the benefit of little theoretical or empirical guidance. In this section, we introduce a theoretical framework to understand the public disclosure of individual ballot records as a special case of election reporting.
Framework for Vote Revelation
To understand the debate about ballot records, it is helpful to conceive of election reporting as a query of an election database. The building blocks of an election database include a table of individual ballot records and a table of individual voter registration records. The ballot records indicate each choice made in each contest by each voter in any election, while the voter registration records enumerate which registrants cast a ballot in the same election. Importantly, the ballot records are anonymous, i.e., no names are attached to the actual ballot, but voter records are not. In this framework, the empirical concern raised by public officials is that a vote choice recorded in a table of election results will be linked to an individual record in a table of voters. We call this vote revelation.
The Venn diagram in Figure 1 illustrates how vote revelation is possible. Both private ballot records and public voter records share overlapping information. At the intersection of the Venn diagram are a voter’s precinct, ballot style, and vote method. Each voter is assigned to a precinct to vote. Further, each voter is assigned a ballot style. We use the term ballot style to refer to the set of contests printed on each voter’s ballot, which differ by jurisdiction (See Section B.1). Finally, vote method indicates whether the voter voted in-person, by mail, or by provisional ballot. Following [47], we refer to these shared variables as quasi-identifiers. They are quasi-identifiers, rather than personal identifiers, in the sense that they generally do not uniquely link a voter with a vote. Instead, they narrow down which voter might have cast which vote.
Figure 1: Schematic of Potential Privacy Violation by Vote Revelation. A Venn diagram of how quasi-identifiers could link vote choice to personally identifying information.
As we describe in detail in the next subsection, quasi-identifiers can lead to vote revelation in specific, well-defined instances by linking the ballot and voter records in Figure 1. The quasi-identifiers are available in voter records because they define where voters vote (a precinct), which contests they vote in (a ballot style), and how they participate (a vote method). Similarly, the quasi-identifiers are available in ballot records because they define the reporting units for aggregating election results. For example, in order to report election results by precinct, election officials would need to know the precinct in which each ballot was cast.
This setup thus mirrors common studies in the privacy literature, where scholars have shown how an analyst might triangulate multiple data sets with overlapping information to gain information not intended to be shared. For example, a typical privacy attack might integrate the description of demographic characteristics of an anonymous medical patient with a database of names and demographics to reveal the name of that patient [[, see, e.g.,]]ohm2009broken,sweeney2017re, yoo2018risks. Here, a voter’s quasi-identifiers—their precinct, ballot style, and vote method—could link a vote choice recorded in a table of election results to an individual record in a table of voters. The privacy literature would describe this as re-identification [[]217]wood2018differential. We prefer the term revelation to emphasize the fact that ballots are anonymous.
Our conception of vote revelation is meant to be responsive to the particular empirical concerns of public officials discussed above. If public officials fear that releasing individual ballot records will reveal vote choices, it is useful to understand the extent to which that concern is realized. However, our conception of vote revelation is neither a sufficient nor necessary condition for determining whether any method of election reporting violates the legal requirement for a secret ballot. It is not a necessary condition because, in general, there is no singular, well-defined secret ballot. Instead, the secret ballot is a state right, enshrined in state constitutions and state statutes in different ways. For example, some state constitutions guarantee “absolute” secrecy and some simply guarantee that voting “shall be by ballot” [25]. Similarly, some permit voters to waive the right and some do not [25]. Given the variation in text and other constitutional commitments, courts—even if not election officials—may also consider other dimensions of the secret ballot, beyond the extent of revelation, such as the harm of revelation or risk of revelation. Vote revelation is also not a sufficient condition for the secret ballot, for the simple reason that public officials still need to supply their own threshold to distinguish some vote revelation from too-much vote revelation. Nonetheless, we argue that our concept of vote revelation is an important yardstick by which to measure the secret ballot.
Mechanism for Revelation
We distinguish between two different types of vote revelation. A vote can either be revealed to (1) the public as a whole or (2) to specific local citizens who are able to supplement the election results and the voter file with knowledge of their own vote (or neighbors’ votes). We term the former public vote revelation in contrast to the latter local vote revelation. Both types of revelation focus on a reporting unit for election results defined by some combination quasi-identifiers, either precinct, ballot style, or vote method.
Public vs. Local Revelations
Figure 2 offers an illustrative example of how we define public and local revelation. Suppose that election results are reported by precinct, ballot style, and vote method. Further, suppose that 30 individuals cast votes in a particular precinct using a particular ballot style, although some individuals vote in person while others vote by mail. Finally, suppose there are just two contests, each with three options. In the contest for President, individuals can vote for Trump, Biden, or abstain. In the second contest on the same ballot, individuals can support, oppose, or abstain from a local tax referendum. In the election administration literature, such abstensions are known as undervoting.
Figure 2: How Unanimous Election Results Reveal Votes. An illustrative example of an election reported at the precinct $\times$ ballot style $\times$ vote method. The example depicts 30 voters in a particular precinct with a ballot style. Only the 10 in-person voters’ votes for President highlighted in yellow can be publicly revealed to everyone. The 19 Trump voters who voted by mail can be locally revealed to the sole Biden voter who voted by mail.
The first row of Figure 2 reports that all 10 in-person voters supported Donald Trump for president. The in-person voters are thus vulnerable to public vote revelation as we defined it above. An analyst could find the names of all ten voters who are registered in precinct A, assigned to ballot style X, and recorded as voting in-person from the public voter file, and know with confidence that they voted for the Donald Trump. We call this public revelation because any analyst anywhere who has access to this public data can link the Trump voters. Importantly, the example shows that there are public revelations of this sort only if all of the voters in a single reporting unit are unanimous. In the example, the 20 mail voters are not vulnerable to public revelation since they were not unanimous (19 Trump voters and 1 Biden voter) —it would not be possible to identify which of the 20 mail voters supported Biden. The example also shows that vote revelation is contest-specific: While the 10 in-person voters would have their presidential vote revealed, the same individuals would not have their referendum vote revealed, because four voters abstained. For the purposes of vote revelation, undervoting is effectively voting for a second candidate.
Given that public revelation relies on unanimity, the privacy literature refers to the mechanism for such revelation as a homogeneity attack [47]. More formally, the literature considers whether or not a dataset is vulnerable to a homogeneity attack based on the concept of $\ell$-diversity. A dataset is considered $\ell$-diverse if there are at least $\ell$ distinct “well-represented” values of a sensitive value in a block of data with a specific quasi-identifier, such as vote method, precinct, or ballot style. [^2] This contrasts with the older notion of $k$-anonymity [62], which holds that a dataset is $k$-anonymous if there are at least $k-1$ individuals with the same quasi-identifier. Even if a dataset is $k$-anonymous, if all $k-1$ voters in a single geography vote unanimously (in other words, $\ell=1$ with only one value), their sensitive attribute is revealed. See our discussion on probabilistic voting for further discussion of the definition of “well-represented.”
Votes that cannot be revealed publicly can still be revealed locally. Local revelation can occur if there is private information available about some individual vote choices in the reporting unit. Local revelation therefore relies on three datasets instead of the two depicted in the Venn diagram of Figure 1: the election results, the voter file, and the voter’s knowledge of their own vote. In Figure 2, the single Biden voter who cast a mail ballot could, upon hearing the 19-1 result in her reporting unit, deduce with certainty that 19 of her neighbors voted for Trump. In other words, she has linked each Trump vote to her neighbors, based on her own knowledge that she voted for Biden. However, each Trump vote would only be revealed to her, not the public as a whole. The extent of local revelation depends on the extent of private information. For example, among mail voters, 17 voted yes on the referendum, 2 voted no, and 1 abstained. If the 3 no-or-abstain voters share their vote choice with each other, they would reveal, but only to themselves, the tax referendum votes of the 17 other voters in their reporting unit. We propose to denote the number of such collaborators required for local revelation by $\alpha$. In our framework, a vote is subject to local revelation at parameter $\alpha$ if the voters in a reporting unit, excluding the $\alpha$ collaborators, are otherwise unanimous.
In formal privacy terminology, the concept of local revelation makes clear that datasets which are $\ell$-diverse are not necessarily immune from vote revelation [14]. Instead, a person who voted themselves can use their own private knowledge to back out how others voted. For this reason, the mechanism for what we call local vote revelation is known as a background knowledge attack [47].
Reporting Units vs. Individual Records
Given the mechanism of vote revelation, releasing individual ballot records is simply a special case of election reporting. That is because (public) vote revelation requires unanimity among voters with the same quasi-identifiers. Consider a precinct where all voters support the same candidate in a given contest. In that case, election results reported at the precinct-level would be unanimous and thus could lead to public vote revelation. So would the release of individual ballot records from that precinct, since ballot records include a quasi-identifier for each precinct and thus can simply be aggregated to the precinct-level.
One concern is that ballots could reveal more information because they show how a single voter voted on all offices. But the example in Figure 2 already showed otherwise. Even if we had individual ballots for the 10 in-person voters, we could not identify how a particular individual voted on the tax referendum even though we could identify that they voted for Trump. In general, individual ballot records have been mischaracterized as posing an exceptional privacy threat. But releasing ballot records can reveal no more votes than would otherwise be revealed by the most granular level of aggregate reporting. [^3] Our claim is specific to our definition of vote revelation. Our definition of vote revelation does not include so-called negative revelation: the ruling out of choices that a voter may have made. Negative revelation is often found in granular election results where it is not uncommon to find zero votes cast for particular minor-party candidates in a reporting unit ruling out the possibility that anyone known to have voted in that reporting unit support that candidate. To be clear, cast vote records do facilitate more negative revelation than aggregate election results. Consider the same two contests as Figure 2, the President and a yes/no referendum. Suppose a vote buyer asks one voter to vote [Trump; Yes] and another to vote [Biden; No]. However, if the first voter voted [Trump; No], and the second voted [Biden; Yes], the cast vote records would reveal that both voters had reneged, while aggregate results would not. We thank Samuel Baltz for providing this example.
It is important to note that there are two different types of individual ballot records, with subtly different implications for public vote revelation. However, the two records have often been confused in the public debate.
Figure 3 compares a cast vote record (left) with a ballot image (right). A cast vote record is a machine-readable enumeration of a voter’s choice across some or all contests on their ballot [65, 45, 46]. Cast vote records are generally stored in familiar data formats such as XML, JSON, or spreadsheets. A ballot image is a digitally recorded scanned image of each ballot or ballot page. Vote scanners can produce both types of individual records.
From a privacy perspective, the relevant difference is that cast vote records only contain a voter’s contest choices, while a ballot image would capture all marks or notations made on the ballot, including those that do not indicate a choice for a particular candidate. For this reason, voters are often prohibited from adding identifying marks to their ballot to identify themselves (See, e.g., Minn. Stat. §204C.22(13)), and election officials attempt to redact ballot images with identifying marks [[]also see Figure B2 for a redacted mark]Colorado2018. Nonetheless, ballot images are more vulnerable to vote revelation than cast records because deterrence and redaction may be ineffective or incomplete. Further, as we describe below, our interest is not in self-revelation, but rather the revelation of vote choices that were not intended to be revealed to anyone. For this reason, we study what the choices themselves can reveal. Those choices are contained in the cast voter records and, for the remainder of this article, we focus specifically those records.
Figure 3: Two Types of Individual Ballot Records. The right image is an example ballot image from Dodge County, Wisconsin. The county publicly releases ballot images from recent elections. The left schematic shows what a cast vote record of this particular ballot would look like to the analyst, using a hypothetical ballot and precinct identifier.
Statistical Model for Vote Revelation
A simple statistical model summarizes the mechanics of vote revelation. Consider a single contest in a reporting unit with $N$ voters. Let the number of publicly revealed votes in that unit be denoted by $R$. In the simplest case, assume there are only two candidates, and no voter abstains from voting. The a priori probability that a randomly selected voter in the reporting unit supports the leading candidate is $w_{1}.$ The probability that all $N$ voters vote for the leading candidate is $w_{1}$ to the $N$th power, $(w_{1})^{N}.$ The probability that all of the voters instead support the trailing candidate is $(1-w_{1})^{N}.$ Therefore, the expected number of revealed votes is
$\mathrm{E}[R]=N\left(w_{1}^{N}+(1-w_{1})^{N}\right),$ (1)
or the number of votes in the reporting unit multiplied by the probability that every voter in the reporting unit makes the same choice.
Based on the simple model, Figure 4 illustrates how expected vote revelation relates to the size of the reporting unit. It considers the typical uncompetitive contest, in which the probability that a randomly selected voter supports the leading candidate is $w_{1}$ = 0.7. When there is a single voter in a reporting unit, that voter’s choice is necessarily revealed, so the expected number of revealed votes is one. The expected number of revealed votes briefly increases when there are two voters in a reporting unit because, in an uncompetitive contest, it is more likely than not that the two voters support the same candidate. After that, expected revelations decrease rapidly as $N$ grows. In this example, expected revelations drop below 0.01 after $N=22$, and drop below 0.001 after $N=29.$
Figure 4: Modelling Expected Revelations. Public vote revelations as a function of the size of a reporting unit. This example depicts a two-candidate contest without abstention where the probability of voting for the leading candidate is 0.7.
Appendix A shows how the simple model can be extended to multi-candidate contests where some voters abstain and can quantify expected local revelation too. In general, the expected degree of revelation is lowest when the abstention rate is highest and candidate support is evenly split among the candidates. In this situation, the number of expected vote revelations approaches zero when as few as 10 votes are tallied in the reporting unit.
Together, the model, figure, and Appendix highlight how vote revelation depends on both administrative choices and individual voting behavior. The election administrator can decrease the number of expected revelations by drawing larger precincts. Relatedly, mapdrawers can decrease expected revelations by drawing political districts that split fewer administrative precincts, and thus also increase the size of the reporting unit. Nonetheless, voters in contests where preferences are clustered, lopsided, or are expected to produce few abstentions will always be more vulnerable to revelation.
Not Vote Revelation
Before proceeding, we clarify the scope of our definition of vote revelation. We set aside three related privacy concerns: about probabilistic updates of vote choices; additional quasi-identifiers on ballots such as time-stamps; and what we term vote authentication. In general, we set these concerns aside because they are not the focus of election officials who are grappling with whether to release cast vote records, as discussed above.
Probabilistic Posterior Updates
First, we do not consider a highly predictive estimate of a voter’s choice to be vote revelation. Our approach is thus distinct from some definitions of statistical privacy, which cast the leakage of private information in Bayesian statistical learning terms. In the alternative approach, a disclosure of data that changes the analyst’s posterior belief about an attribute’s value is treated as a successful privacy attack [37, 17, 41]. [^4] In these cases, it is typical to suppose that the attacker already holds accurate prior knowledge of the joint distribution of the sensitive items (votes) and the quasi-identifiers where the joint distribution is defined over a hyper-population of possible databases. If the information presented in the database does not substantially update the attacker’s beliefs about the values of the sensitive items (vote choices) of any given voter from their prior belief (based on the individual’s quasi-identifier attributes), then it does not leak privacy. But that approach fits uneasily in the U.S. election setting, where campaigns are flush with individually-identifiable information about voters [35, 57]. In the current political environment, experts already often have strong prior beliefs about most voters’ vote choices, without access to ballot records. Campaigns rely on a combination of precinct-level election results, party registration, and campaign contributions. As a result, short of revelation, there is arguably little left to learn about any specific individual’s vote choices from ballot records.
Other Quasi-identifiers
Second, we do not consider the possibility of vote revelation from quasi-identifiers other than precinct, ballot style, and vote method. For example, election officials have voiced concern that including time-stamps on ballots records indicating exactly when each ballot was cast (or in what order the ballots were cast) could be used to reveal votes (See, e.g., Honey v. Lycoming County). [^5] In fact, at least one voting system is known to have inadvertently disclosed the order in which in-person ballots were cast [18]. Assuming the order in which individuals voted is observable or recorded in a poll book, the ability to sort the ballots in the order in which they were cast is fully revealing of which voters cast which ballot. We set aside such concerns arising from other quasi-identifiers because the solution to these concerns is about election administration, not the reporting of election results. To continue the example, election officials need not provide a time-stamp of a voter’s interaction with a ballot in a publicly available cast vote record. In contrast, if election officials are required to report precinct-level results, they must necessarily associate the voter’s precinct with each ballot in order to administer the election. Similarly, information about the ballot style is inherent to each ballot and cannot be redacted.666 The cast vote records produced by some voting systems do not connect vote choices made by voters across a multiple-page ballot. In this case, ballot style may not be entirely revealed by the cast vote records and, thus, ballot style information is not inherent to cast vote records in those cases. In the limit, a cast vote record could report vote choices by contest and not link voter choices across any pair of contests for any voter. Such a cast vote records database would be informationally equivalent to a database that contained only candidate vote totals. Election officials have also voiced concern about ballots in different languages, but we are not presently aware of ballot language being recorded in a voter file.
Authentication
Finally, our definition of vote revelation does not consider the possibility of voters using cast vote records to voluntarily reveal their vote. Prior work has warned that releasing ballot images could allow voters to authenticate their votes by adding identifiable stray marks, which could be captured in ballot images [2]. That concern is also reflected in state law. For example, as mentioned above, Minnesota requires that a ballot not be intentionally marked with “distinguishing characteristics” that render it identifiable (Minn. Stat. §204C.22(13) (2022)). More recent work focuses on how releasing cast vote records in ranked-choice elections could similarly allow each voter to use their complete ranking ordering of the candidates as a verifiable public signal that would nearly always be unique to them [67].
Nonetheless, we see vote revelation as a more pressing concern than vote authentication in considering the public disclosure of individual ballots. For one, vote-buying and vote-selling already violates federal law [32]. Further, vote authentication is often permissible as protected First Amendment speech, at least in some jurisdictions (See, e.g., Rideout v. Gardner, 838 F.3d 65 (1st Cir. 2016)). For example, many states permit voters to share a picture of their marked ballot, known as a “ballot-selfie” [52]. As a result, vote revelation concerns are more relevant than vote authentication concerns to the disclosure of cast vote records.
Data and Case Selection
We use Maricopa County, Arizona, as a case study to quantify the extent of vote revelation in the reporting of election results. [^7] Arizona features both a Maricopa County and a town of Maricopa, which is in Pinal County. All references to Maricopa are to the county.
Case Selection
Following the 2020 presidential election, Maricopa County became the most prominent site in the battle over public access to ballots, and perhaps over the legitimacy of U.S. elections. Maricopa responded to the attention by adopting a transparency approach and releasing its cast vote records.
Importantly, Maricopa’s cast vote records include the necessary information on both quasi-identifiers and vote choice to allow us to reconstruct election results at any common level of aggregation. For each contest on each ballot, Maricopa’s cast vote records provide vote choice, including undervotes or overvotes, vote method (in-person, early, provisional), precinct, and ballot style. [^8] Since 2021, Maricopa has redacted the date stamp fields from their cast vote record.
The county is a useful case study also because it is closely divided by partisanship, populous with over 2 million voters in the 2020 election, and features a range of geographies from urban, suburban, and rural areas. Finally, Maricopa’s 2020 general election also contains hundreds of contested elections. Figure C5 shows a summary of the contests we examine.
Data Processing
Table 1: Summary Statistics. Counts and coverage of the Maricopa County data we focus on in this paper. Each row shows, for each type of reporting unit, the distribution of the voter size of each unit. The column “unique unit” indicates the number of unique units that exist. Columns labeled by percentages indicate the percentile of the distribution of the number of voters in each unit.
Quantiles and mean of voters in unit
Reporting Unit Unique unit 1% 5% 10% 50% Mean
Precinct 743 33 602 1,090 2,600
Ballot Style & 381 1 7 41 1,897 5,484.4 Precinct $\times$ Vote Method 2,195 1 3 5 208 952.0 Ballot equivalent 4,397 1 1 1 11 475.2     … only mail method 1,742 1 2 4 257 1,099.6     … only in-person method 1,389 1 1 1 87 120.9     … only provisional method 1,320 1 1 1 3 4.7 Cast vote record 2,089,563 We use Ballot to refer to Precinct $\times$ Ballot Style $\times$ Vote Method units, as we describe in the main text.
Table 4.2 summarizes our collection of cast vote records for the November 2020 general election. The bottom row of Table 4.2 shows that we can reconstruct all 2,089,563 valid ballots tabulated in the November 2020 election from the cast vote records. The remaining rows detail how cast vote records can be used to reconstruct election results at various reporting levels. For example, the top row shows that Maricopa County used 743 unique precincts for the election. The median precinct included 2,600 voters, while the precinct at the 1st percentile of the distribution contained only 33 voters. The second row shows that there were 381 unique ballot styles, with at least 1 percent of the 381 ballot styles containing only 1 voter, and the median ballot style containing 1,897 voters. [^9] A large contributor to ballot styles with extremely small populations in Arizona is the provision of a federal-only ballot, in which voters without sufficient state IDs are only eligible for the federal offices in the general election. These are treated as a separate ballot style in each precinct. The most granular reporting unit is precinct $\times$ ballot style $\times$ vote method, which we have established is equivalent to releasing individual ballots. With three types of voting methods in the cast vote record data, there are 4,397 such combinations of precinct-style-method with at least one vote. The three rows with “…” further subset style-method reporting units by each of the three vote methods. Reporting units for provisional ballots have by far the fewest number of voters.