### 09 Apr 2024 v3

# The Still Secret Ballot:

Abstract
After an election, should officials release an electronic record of each ballot? The release of ballots could bolster the legitimacy of the result. But it may also facilitate vote revelation, where an analyst unravels the secret ballot by uniquely linking votes on an anonymous ballot to the voter’s name and address in the public voter file. We first provide a theoretical model of how vote revelation could occur under various election-reporting regimes. Perhaps counterintuitively, releasing ballot records is no more revelatory than the typical practice of releasing aggregate vote tallies by precinct and method. We then present the first empirical evaluation of vote revelation, using the 2020 election in Maricopa County, Arizona, as a case study. For 99.8% of voters, the release of ballot records led to no revelation of any vote choice. We conclude the ballot can be both public and still as secret as it is under typical reporting practices.
Introduction
Democracies face a difficult tradeoff in election reporting. After holding an election, election officials must report results at a level of granularity sufficient to bolster the legitimacy of the vote count while avoiding revealing individual vote choices. Where to strike the balance between transparency and privacy is an increasingly pressing question: A growing swath of voters who do not believe their votes are counted accurately are seeking “citizen audits” of elections [61, 56, 10].
States and localities vary in how they report election results. North Carolina offers a particularly illuminating example of the transparency-privacy tradeoff. The State Board of Elections is obligated to both report granular election results by precinct and vote method and protect the state constitutional right to a secret ballot (N.C. Gen. Stat. §163-132.5G). To strike that balance, the board adds or subtracts an undisclosed random number to some reported election results. The board is concerned with what we term vote revelation: an analyst uniquely linking the anonymous vote choices reflected in election results with the public record of voters’ name and addresses. However, as the board acknowledges, the random noise means that a candidate who appears to have won the most votes in the published noised election results may not be the one who was the actual certified winner. [^1] The state website explains that noising may at times have the appearance “of changing the outcome of very close contests.” See https://perma.cc/URR8-Z5YQ and https://perma.cc/F646-EG4Y.
The downside of adding noise to election results is that it makes the results less accurate and possibly misleading [37]. Such privacy-protecting manipulations can also impede the application of election forensics to detect genuine electoral manipulation or malfunction [[, see, e.g.,]]mebane2008election, cantu2014identifying, callen2015institutional, cottrell2018exploration. In short, granular election results are important in order for elections to be accepted as legitimate. At the same time, voter privacy contributes to the legitimacy of the result too. In fact, the secret ballot was adopted to reduce corruption and voter intimidation [43, 48].
This study presents the first empirical estimates of privacy loss under different election reporting regimes, including the public release of anonymous ballots. We use the complete set of individual-level ballot records, known as cast vote records, in Maricopa County, Arizona, as the primary case study. The dataset allows us to compute the extent of potential privacy violations under various reporting regimes, holding voting behavior constant. We estimate that the public release of cast vote records could reveal about 0.0003% of votes for President in the November 2020 election, among the over 2 million voters we consider. Moreover, for more than 99.8% of voters, no revelation of any vote choice in any contest was possible across the 60-70 contests that they voted on.
We conclude that the ballot can still be considered secret even if cast vote records are subsequently made public, for two reasons we discuss in detail. First, the fraction of potential votes revealed is small enough that privacy concerns can be addressed without withholding all cast vote records. Second, the release of cast vote records is no more revelatory than the current practice of reporting aggregate vote tallies by precinct and vote method. The ballot is thus still secret in the sense that the release of cast vote records does not substantially change the privacy risks that already exist when jurisdictions release granular aggregate election results.
Our argument thus differs from the privacy-centric view put forth in academia and the courts. For example, [2] argue against releasing images of each ballot cast because of possible risks to the secret ballot. Further, the top state court in New York ruled that anonymous individual ballot records were not public record in part because they might reveal vote choices. The Governor of Arizona recently vetoed a bipartisan bill to release cast vote records for a similar reason [36]. But neither [2] nor the New York court nor the governor of Arizona estimated the extent of vote revelation.
Our study offers three contributions. First, we propose a way for election administration to incorporate concepts from the burgeoning field of privacy. In the last two decades, a movement largely led by computer scientists has shown how ostensibly private datasets can have their privacy undone by data triangulation [1, 62, 23, 68]. But importing the dominant approach in data privacy to the electoral context is not straightforward. Election data has several distinct features compared to the commercial or medical datasets commonly used in privacy studies: ballots themselves do not contain personally identifying information, the certified result must provide finality, and votes should be anonymous not just to the public, but also to election officials.
Second, using our framework, we provide the first empirical assessment of the privacy costs of transparent election results. Although a limited set of academic studies and assessments by expert commissions has already pointed to the potential for vote revelation from releasing cast vote records, those studies do not go on to measure the extent of the problem in an actual election [2, 50, 8], or, if they do, only measure a special case of vote revelation [13]. Our study clarifies that the reporting of election results should only lead to what we term vote revelation when there is a unanimous result in a reporting unit (for example, in a precinct of 10 voters, all 10 of them vote for the Republican candidate). Further, it shows how the release of cast vote records is no more revelatory than the release of election results aggregated by the partially identifying information (what we call quasi-identifiers) available in cast vote records.
Third, and finally, the extent of revelation is directly relevant to current litigation about whether cast vote records should be public record for transparency or kept confidential for privacy reasons. The extent of revelation is also pertinent to policy decisions about how to best strike the difficult balance between transparency and privacy in election reporting. We estimate that less than 3,500 out of over 2 million Maricopa County voters could have any of their votes revealed. Further, when votes are revealed, they are disproportionately provisionally cast ballots ( which only comprise 0.1 percent of total ballots cast), and are usually votes for candidates who are popular in the area in which the voter lives. These findings help inform whether a particular response—ranging from not releasing any individual ballots to aggregating, suppressing, or adding random noise to results—is proportionate to the potential privacy loss.
Current Challenges in Election Reporting
Across the country, election administrators and state officials alike are currently grappling with how to best report election results. In general, election administrators have typically tallied individual ballots and reported aggregate results. For example, almost every state reported results for the 2016 and 2020 presidential elections results by precinct [6]. But vote tabulation technology has been changing. Most modern voting systems can also preserve anonymous, individual ballot records in addition to tallying aggregate results. The changing technology has made clear the tradeoff between voter privacy and public transparency in election reporting.
The rise of election skepticism has triggered increased interest in individual ballot records. A number of election officials now affirmatively release these ballot records, including the states of Alaska and Maryland, as well as the counties of San Francisco, California, Dane County, Wisconsin, and Leon County, Florida [4]. But many other election officials have prohibited the release of ballot records. For example, a review of recent news reports reveals that election officials in Indiana, Missouri, and South Carolina have all denied requests for ballot records [58, 70, 38].
Election officials have come to different conclusions in part because of an empirical disagreement about the consequences of public disclosure [24]. For example, the Alaska Division of Elections has released individual ballots because “ballots are anonymous and not connected to voters” [7]. Similarly, Dane County has noted that “there [is] nothing linking the data to individual voters.” In contrast, the South Carolina Election Commission has denied public records requests because “the release of [ballot records] would likely lead to [the] identification of voters” [39]. Similarly, the Indiana Secretary of State Election Division determined that disclosure “would identify each voter” [54]. The conflict has come to bear within states too. For example, Arizona’s Democratic Secretary of State recently supported a bill to release individual ballots because it “would help restore trust in [the] election process” [26], but the Democratic governor ultimately vetoed it, arguing it “threatens anonymity” [36].
State courts have also reached conflicting conclusions about the release of ballot records because of the same empirical disagreement. Their legal analysis focuses on the specifics of state public records law. In general, under the typical state public records law, the release of ballots depends on whether public disclosure would violate the state right to a secret ballot. For example, a Pennsylvania court recently held that the release of ballot records does not violate the secret ballot provision in the state’s constitution because disclosure will not reveal voters’ selections (Honey v. Lycoming County, No. 22-cv-115 (Pa. Commw. Ct. Dec. 15, 2022)). In contrast, the New York Court of Appeals, the highest court in the state, has come to the opposite conclusion. (Kosmider v. Whitney, 132 N.E.3d 592 (N.Y. 2019)). There is also active litigation in other states. For example, a Missouri county clerk recently filed suit to seek clarification about the public status of individual ballot records (Petition for Declaratory Judgement, Schoeller v. Huddleston, No. 2231-CC00975 (Greene County, Mo. Cir. Ct. 2022)).
Vote Revelation
The lack of consensus about the release of individual ballot records reflects, at least in part, the lack of empirical research. Election officials and courts alike have been left to navigate the new technological landscape with the benefit of little theoretical or empirical guidance. In this section, we introduce a theoretical framework to understand the public disclosure of individual ballot records as a special case of election reporting.
Framework for Vote Revelation
To understand the debate about ballot records, it is helpful to conceive of election reporting as a query of an election database. The building blocks of an election database include a table of individual ballot records and a table of individual voter registration records. The ballot records indicate each choice made in each contest by each voter in any election, while the voter registration records enumerate which registrants cast a ballot in the same election. Importantly, the ballot records are anonymous, i.e., no names are attached to the actual ballot222 One caveat is that North Carolina marks some ballots—those cast during early voting—with an identifier such that it is “retrievable” (N.C. Gen. Stat. §163-166.45). To be clear, “[t]he ballot number or equivalent identifier shall not be a public record.” Instead, the ballot identifier allows election officials to remove a ballot from counting if it is later found to be fraudulent. While making the ballot retrievable means that it may no longer be anonymous to election officials, it is a criminal offense for anyone “who has access to an official voted ballot or record and knowingly discloses … how an individual has voted that ballot” (N.C. Gen. Stat. §163-165.1(e))., but voter records are not. In this framework, the empirical concern raised by public officials is that a vote choice recorded in a table of election results will be linked to an individual record in a table of voters. We call this vote revelation.
The Venn diagram in Figure 1 illustrates how vote revelation is possible. Both private ballot records and public voter records share overlapping information. At the intersection of the Venn diagram are a voter’s precinct, ballot style, and vote method. Each voter is assigned to a precinct to vote. Further, each voter is assigned a ballot style. We use the term ballot style to refer to the set of contests printed on each voter’s ballot, which differ by jurisdiction (See Section B.1). Finally, vote method indicates whether the voter voted in-person, by mail, or by provisional ballot. Following [47], we refer to these shared variables as quasi-identifiers. They are quasi-identifiers, rather than personal identifiers, in the sense that they generally do not uniquely link a voter with a vote. Instead, they narrow down which voter might have cast which vote.
Figure 1: Schematic of Potential Privacy Violation by Vote Revelation. A Venn diagram of how quasi-identifiers could link vote choice to personally identifying information.
As we describe in detail in the next subsection, quasi-identifiers can lead to vote revelation in specific, well-defined instances by linking the ballot and voter records in Figure 1. The quasi-identifiers are available in voter records because they define where voters vote (a precinct), which contests they vote in (a ballot style), and how they participate (a vote method). Similarly, the quasi-identifiers are available in ballot records because they define the reporting units for aggregating election results. For example, in order to report election results by precinct, election officials would need to know the precinct in which each ballot was cast.
This setup thus mirrors common studies in the privacy literature, where scholars have shown how an analyst might triangulate multiple data sets with overlapping information to gain information not intended to be shared. For example, a typical privacy attack might integrate the description of demographic characteristics of an anonymous medical patient with a database of names and demographics to reveal the name of that patient [[, see, e.g.,]]ohm2009broken,sweeney2017re, yoo2018risks. Here, a voter’s quasi-identifiers—their precinct, ballot style, and vote method—could link a vote choice recorded in a table of election results to an individual record in a table of voters. The privacy literature would describe this as re-identification [[]217]wood2018differential. We prefer the term revelation to emphasize the fact that ballots are anonymous.
Our conception of vote revelation is meant to be responsive to the particular empirical concerns of public officials discussed above. If public officials fear that releasing individual ballot records will reveal vote choices, it is useful to understand the extent to which that concern is realized. However, our conception of vote revelation is neither a sufficient nor necessary condition for determining whether any method of election reporting violates the legal requirement for a secret ballot. It is not a necessary condition because, in general, there is no singular, well-defined secret ballot. Instead, the secret ballot is a state right, enshrined in state constitutions and state statutes in different ways. For example, some state constitutions guarantee “absolute” secrecy and some simply guarantee that voting “shall be by ballot” [25]. Similarly, some permit voters to waive the right and some do not [25]. Given the variation in text and other constitutional commitments, courts—even if not election officials—may also consider other dimensions of the secret ballot, beyond the extent of revelation, such as the harm of revelation or risk of revelation. Vote revelation is also not a sufficient condition for the secret ballot, because public officials still need to supply their own threshold to distinguish some vote revelation from too much vote revelation. Nonetheless, we argue that our concept of vote revelation is an important yardstick by which to measure the secret ballot.
Mechanism for Revelation
We distinguish between two different types of vote revelation. A vote can either be revealed to (1) the public as a whole or (2) to specific local citizens who are able to supplement the election results and the voter file with knowledge of their own vote (or neighbors’ votes). We term the former public vote revelation in contrast to the latter local vote revelation. Both types of revelation focus on a reporting unit for election results defined by some combination of quasi-identifiers, either precinct, ballot style, or vote method.
Public vs. Local Revelations
Figure 2 offers an illustrative example of how we define public and local revelation. Suppose that election results are reported by precinct, ballot style, and vote method. Further, suppose that 30 individuals cast votes in a particular precinct using a particular ballot style, although some individuals vote in person while others vote by mail. Finally, suppose there are just two contests, each with three options. In the contest for President, individuals can vote for Trump, Biden, or abstain. In the second contest on the same ballot, individuals can support, oppose, or abstain from a local tax referendum. In the election administration literature, such abstensions are known as undervoting.
Figure 2: How Unanimous Election Results Reveal Votes. An illustrative example of an election reported at the precinct $\times$ ballot style $\times$ vote method. The example depicts 30 voters in a particular precinct with a ballot style. Only the 10 in-person voters’ votes for President highlighted in yellow can be publicly revealed to everyone. The 19 Trump voters who voted by mail can be locally revealed to the sole Biden voter who voted by mail.
The first row of Figure 2 reports that all 10 in-person voters supported Donald Trump for president. The in-person voters are thus vulnerable to public vote revelation as we defined it above. An analyst could find the names of all ten voters who are registered in precinct A, assigned to ballot style X, and recorded as voting in-person from the public voter file, and know with confidence that they voted for the Donald Trump. We call this public revelation because any analyst anywhere who has access to this public data can link the Trump voters. Importantly, the example shows that there are public revelations of this sort only if all of the voters in a single reporting unit are unanimous. In the example, the 20 mail voters are not vulnerable to public revelation since they were not unanimous (19 Trump voters and 1 Biden voter) —it would not be possible to identify which of the 20 mail voters supported Biden. The example also shows that vote revelation is contest-specific: While the 10 in-person voters would have their presidential vote revealed, the same individuals would not have their referendum vote revealed, because four voters abstained. For the purposes of vote revelation, undervoting is effectively voting for a second candidate.
Given that public revelation relies on unanimity, the privacy literature refers to the mechanism for such revelation as a homogeneity attack [47]. More formally, the literature considers whether or not a dataset is vulnerable to a homogeneity attack based on the concept of $\ell$-diversity. A dataset is considered $\ell$-diverse if there are at least $\ell$ distinct “well-represented” values of a sensitive value in a block of data with a specific quasi-identifier, such as vote method, precinct, or ballot style. [^3] This contrasts with the older notion of $k$-anonymity [62], which holds that a dataset is $k$-anonymous if there are at least $k-1$ individuals with the same quasi-identifier. Even if a dataset is $k$-anonymous, if all $k-1$ voters in a single geography vote unanimously (in other words, $\ell=1$ with only one value), their sensitive attribute is revealed. See our discussion on probabilistic voting for further discussion of the definition of “well-represented.”
Votes that cannot be revealed publicly can still be revealed locally. Local revelation can occur if there is private information available about some individual vote choices in the reporting unit. Local revelation therefore relies on three datasets instead of the two depicted in the Venn diagram of Figure 1: the election results, the voter file, and the voter’s knowledge of their own vote. In Figure 2, the single Biden voter who cast a mail ballot could, upon hearing the 19-1 result in her reporting unit, deduce with certainty that 19 of her neighbors voted for Trump. In other words, she has linked each Trump vote to her neighbors, based on her own knowledge that she voted for Biden. However, each Trump vote would only be revealed to her, not the public as a whole. The extent of local revelation depends on the extent of private information. For example, among mail voters, 17 voted yes on the referendum, 2 voted no, and 1 abstained. If the 3 no-or-abstain voters share their vote choice with each other, they would reveal, but only to themselves, the tax referendum votes of the 17 other voters in their reporting unit. We propose to denote the number of such collaborators required for local revelation by $\alpha$. In our framework, a vote is subject to local revelation at parameter $\alpha$ if the voters in a reporting unit, excluding the $\alpha$ collaborators, are otherwise unanimous.
In formal privacy terminology, the concept of local revelation makes clear that datasets that are $\ell$-diverse are not necessarily immune from vote revelation [14]. Instead, a person who voted themselves can use their own private knowledge to back out how others voted. For this reason, the mechanism for what we call local vote revelation is known as a background knowledge attack [47].
Reporting Units vs. Individual Records
Given the mechanism of vote revelation, releasing individual ballot records is simply a special case of election reporting. That is because (public) vote revelation requires unanimity among voters with the same quasi-identifiers. Consider a precinct where all voters support the same candidate in a given contest. In that case, election results reported at the precinct-level would be unanimous and thus could lead to public vote revelation. So would the release of individual ballot records from that precinct, since ballot records include a quasi-identifier for each precinct and thus can simply be aggregated to the precinct-level.
One concern is that ballots could reveal more information because they show how a single voter voted on all offices. But the example in Figure 2 already showed otherwise. Even if we had individual ballots for the 10 in-person voters, we could not identify how a particular individual voted on the tax referendum even though we could identify that they voted for Trump. In general, individual ballot records have been mischaracterized as posing an exceptional privacy threat. But releasing ballot records can reveal no more votes than would otherwise be revealed by the most granular level of aggregate reporting. [^4] Our claim is specific to our definition of vote revelation. Our definition of vote revelation does not include so-called negative revelation: the ruling out of choices that a voter may have made. Negative revelation is often found in granular election results where it is not uncommon to find zero votes cast for particular minor-party candidates in a reporting unit, ruling out the possibility that anyone known to have voted in that reporting unit supports that candidate. To be clear, cast vote records do facilitate more negative revelation than aggregate election results. Consider the same two contests as Figure 2, the President and a yes/no referendum. Suppose a vote buyer asks one voter to vote [Trump; Yes] and another to vote [Biden; No]. However, if the first voter voted [Trump; No], and the second voted [Biden; Yes], the cast vote records would reveal that both voters had reneged, while aggregate results would not. We thank Samuel Baltz for providing this example.
It is important to note that there are two different types of individual ballot records, with subtly different implications for public vote revelation. However, the two records have often been confused in the public debate.
Figure 3 compares a cast vote record (left) with a ballot image (right). A cast vote record is a machine-readable enumeration of a voter’s choice across some or all contests on their ballot [65, 45, 46]. Cast vote records are generally stored in familiar data formats such as XML, JSON, or spreadsheets. A ballot image is a digitally recorded scanned image of each ballot or ballot page. Vote scanners can produce both types of individual records.
From a privacy perspective, the relevant difference is that cast vote records only contain a voter’s contest choices, while a ballot image would capture all marks or notations made on the ballot, including those that do not indicate a choice for a particular candidate. For this reason, voters are often prohibited from adding identifying marks to their ballot to identify themselves (See, e.g., Minn. Stat. §204C.22(13)), and election officials attempt to redact ballot images with identifying marks [[]also see Figure B2 for a redacted mark]Colorado2018. Nonetheless, ballot images are more vulnerable to vote revelation than cast records because deterrence and redaction may be ineffective or incomplete. Further, as we describe below, our interest is not in self-revelation, but rather the revelation of vote choices that were not intended to be revealed to anyone. For this reason, we study what the choices themselves can reveal. Those choices are contained in the cast voter records and, for the remainder of this article, we focus specifically on those records.
Figure 3: Two Types of Individual Ballot Records. The right image is an example ballot image from Dodge County, Wisconsin. The county publicly releases ballot images from recent elections. The left schematic shows what a cast vote record of this particular ballot would look like to the analyst, using a hypothetical ballot and precinct identifier.
Statistical Model for Vote Revelation
A simple statistical model summarizes the mechanics of vote revelation. Consider a single contest in a reporting unit with $N$ voters. Let the number of publicly revealed votes in that unit be denoted by $R$. In the simplest case, assume there are only two candidates, and no voter abstains from voting. The a priori probability that a randomly selected voter in the reporting unit supports the leading candidate is $w_{1}.$ The probability that all $N$ voters vote for the leading candidate is $w_{1}$ to the $N$th power, $(w_{1})^{N}.$ The probability that all of the voters instead support the trailing candidate is $(1-w_{1})^{N}.$ Therefore, the expected number of revealed votes is
$\mathrm{E}[R]=N\left(w_{1}^{N}+(1-w_{1})^{N}\right),$ (1)
or the number of votes in the reporting unit multiplied by the probability that every voter in the reporting unit makes the same choice.
Based on the simple model, Figure 4 illustrates how expected vote revelation relates to the size of the reporting unit. It considers the typical uncompetitive contest, in which the probability that a randomly selected voter supports the leading candidate is $w_{1}$ = 0.7. When there is a single voter in a reporting unit, that voter’s choice is necessarily revealed, so the expected number of revealed votes is one. The expected number of revealed votes briefly increases when there are two voters in a reporting unit because, in an uncompetitive contest, it is more likely than not that the two voters support the same candidate. After that, expected revelations decrease rapidly as $N$ grows. In this example, expected revelations drop below 0.01 after $N=22$, and drop below 0.001 after $N=29.$
Figure 4: Modelling Expected Revelations. Public vote revelations as a function of the size of a reporting unit. This example depicts a two-candidate contest without abstention where the probability of voting for the leading candidate is 0.7.
Appendix A shows how the simple model can be extended to multi-candidate contests where some voters abstain and can quantify expected local revelation too. In general, the expected degree of revelation is lowest when the abstention rate is highest and candidate support is evenly split among the candidates. In this situation, the number of expected vote revelations approaches zero when as few as 10 votes are tallied in the reporting unit.
Together, the model, figure, and Appendix highlight how vote revelation depends on both administrative choices and individual voting behavior. The election administrator can decrease the number of expected revelations by drawing larger precincts. Relatedly, mapdrawers can decrease expected revelations by drawing political districts that split fewer administrative precincts, and thus also increase the size of the reporting unit. Nonetheless, voters in contests where preferences are clustered, lopsided, or are expected to produce few abstentions will always be more vulnerable to revelation.
Not Vote Revelation
Before proceeding, we clarify the scope of our definition of vote revelation. We set aside three related privacy concerns: about probabilistic updates of vote choices; additional quasi-identifiers on ballots such as time-stamps; and what we term vote authentication. In general, we set these concerns aside because they are not the focus of election officials who are grappling with whether to release cast vote records, as discussed above.
Probabilistic Posterior Updates
First, we do not consider a highly predictive estimate of a voter’s choice to be vote revelation. Our approach is thus distinct from some definitions of statistical privacy, which cast the leakage of private information in Bayesian statistical learning terms. In the alternative approach, a disclosure of data that changes the analyst’s posterior belief about an attribute’s value is treated as a successful privacy attack [37, 17, 41]. [^5] In these cases, it is typical to suppose that the attacker already holds accurate prior knowledge of the joint distribution of the sensitive items (votes) and the quasi-identifiers where the joint distribution is defined over a hyper-population of possible databases. If the information presented in the database does not substantially update the attacker’s beliefs about the values of the sensitive items (vote choices) of any given voter from their prior belief (based on the individual’s quasi-identifier attributes), then it does not leak privacy. But that approach fits uneasily in the U.S. election setting, where campaigns are flush with individually-identifiable information about voters [35, 57]. In the current political environment, experts already often have strong prior beliefs about most voters’ vote choices, without access to ballot records. Campaigns rely on a combination of precinct-level election results, party registration, and campaign contributions. As a result, short of revelation, there is arguably little left to learn about any specific individual’s vote choices from ballot records.
Other Quasi-identifiers
Second, we do not consider the possibility of vote revelation from quasi-identifiers other than precinct, ballot style, and vote method. For example, election officials have voiced concern that including time-stamps on ballots records indicating exactly when each ballot was cast (or in what order the ballots were cast) could be used to reveal votes (See, e.g., Honey v. Lycoming County). [^6] In fact, at least one voting system is known to have inadvertently disclosed the order in which in-person ballots were cast [18]. Assuming the order in which individuals voted is observable or recorded in a poll book, the ability to sort the ballots in the order in which they were cast is fully revealing of which voters cast which ballot. We set aside such concerns arising from other quasi-identifiers because the solution to these concerns is about election administration, not the reporting of election results. To continue the example, election officials need not provide a time-stamp of a voter’s interaction with a ballot in a publicly available cast vote record. In contrast, if election officials are required to report precinct-level results, they must necessarily associate the voter’s precinct with each ballot in order to administer the election. Similarly, information about the ballot style is inherent to each ballot and cannot be redacted.777 The cast vote records produced by some voting systems do not connect vote choices made by voters across a multiple-page ballot. In this case, ballot style may not be entirely revealed by the cast vote records and, thus, ballot style information is not inherent to cast vote records in those cases. In the limit, a cast vote record could report vote choices by contest and not link voter choices across any pair of contests for any voter. Such a cast vote records database would be informationally equivalent to a database that contained only candidate vote totals. Election officials have also voiced concern about ballots in different languages, but we are not presently aware of ballot language being recorded in a voter file.
Authentication
Finally, our definition of vote revelation does not consider the possibility of voters using cast vote records to voluntarily reveal their vote. Prior work has warned that releasing ballot images could allow voters to authenticate their votes by adding identifiable stray marks, which could be captured in ballot images [2]. That concern is also reflected in state law. For example, as mentioned above, Minnesota requires that a ballot not be intentionally marked with “distinguishing characteristics” that render it identifiable (Minn. Stat. §204C.22(13) (2022)). More recent work focuses on how releasing cast vote records in ranked-choice elections could similarly allow each voter to use their complete ranking ordering of the candidates as a verifiable public signal that would nearly always be unique to them [67].
Nonetheless, we see vote revelation as a more pressing concern than vote authentication in considering the public disclosure of individual ballots. For one, vote-buying and vote-selling already violates federal law [32]. Further, vote authentication is often permissible as protected First Amendment speech, at least in some jurisdictions (See, e.g., Rideout v. Gardner, 838 F.3d 65 (1st Cir. 2016)). For example, many states permit voters to share a picture of their marked ballot, known as a “ballot-selfie” [52]. As a result, vote revelation concerns are more relevant than vote authentication concerns to the disclosure of cast vote records.
Data and Case Selection
We use Maricopa County, Arizona, as a case study to quantify the extent of vote revelation in the reporting of election results. [^8] Arizona features both a Maricopa County and a town of Maricopa, which is in Pinal County. All references to Maricopa are to the county.
Case Selection
Following the 2020 presidential election, Maricopa County became the most prominent site in the battle over public access to ballots, and perhaps over the legitimacy of U.S. elections. Maricopa responded to the attention by adopting a transparency approach and releasing its cast vote records.
Importantly, Maricopa’s cast vote records include the necessary information on both quasi-identifiers and vote choice to allow us to reconstruct election results at any common level of aggregation. For each contest on each ballot, Maricopa’s cast vote records provide vote choice, including undervotes or overvotes, vote method (in-person, early, provisional), precinct, and ballot style. [^9] Since 2021, Maricopa has redacted the date stamp fields from their cast vote record.
The county is a useful case study also because it is closely divided by partisanship, populous with over 2 million voters in the 2020 election, and features a range of geographies from urban, suburban, and rural areas. Finally, Maricopa’s 2020 general election also contains hundreds of contested elections. Figure C5 shows a summary of the contests we examine.
Data Processing
Table 1: Summary Statistics. Counts and coverage of the Maricopa County data we focus on in this paper. Each row shows, for each type of reporting unit, the distribution of the voter size of each unit. The column “unique unit” indicates the number of unique units that exist. Columns labeled by percentages indicate the percentile of the distribution of the number of voters in each unit.
Quantiles and mean of voters in unit
Reporting Unit Unique unit 1% 5% 10% 50% Mean
Precinct 743 33 602 1,090 2,600 2,812.3
Ballot Style 381 1 7 41 1,897 5,484.4
Precinct $\times$ Vote Method 2,195 1 3 5 208 952.0
Ballot equivalent 4,397 1 1 1 11 475.2
    … only mail method 1,742 1 2 4 257 1,099.6
    … only in-person method 1,389 1 1 1 87 120.9
    … only provisional method 1,320 1 1 1 3 4.7
Cast vote record 2,089,563
We use Ballot to refer to Precinct $\times$ Ballot Style $\times$ Vote Method units, as we describe in the main text.
Table 1 summarizes our collection of cast vote records for the November 2020 general election. The bottom row of Table 1 shows that we can reconstruct all 2,089,563 valid ballots tabulated in the November 2020 election from the cast vote records. The remaining rows detail how cast vote records can be used to reconstruct election results at various reporting levels. For example, the top row shows that Maricopa County used 743 unique precincts for the election. The median precinct included 2,600 voters, while the precinct at the 1st percentile of the distribution contained only 33 voters. The second row shows that there were 381 unique ballot styles, with at least 1 percent of the 381 ballot styles containing only 1 voter, and the median ballot style containing 1,897 voters. [^10] A large contributor to ballot styles with extremely small populations in Arizona is the provision of a federal-only ballot, in which voters without sufficient state IDs are only eligible for the federal offices in the general election. These are treated as a separate ballot style in each precinct. The most granular reporting unit is precinct $\times$ ballot style $\times$ vote method, which we have established is equivalent to releasing individual ballots. With three types of voting methods in the cast vote record data, there are 4,397 such combinations of precinct-style-method with at least one vote. The three rows with “…” further subset style-method reporting units by each of the three vote methods. Reporting units for provisional ballots have by far the fewest number of voters.
Voter File Data
The possibility of vote revelation depends on the completeness of the voter file. As explained above, a voter file includes both personally identifying information (e.g, voters’ names and addresses) and quasi-identifiers (e.g., precinct and vote method). The difficulty is that a voter file is dynamic. As a result, a voter who is currently assigned to a particular precinct may not have been assigned to that precinct during any previous election. In fact, it is well known in the election administration community that public voter files suffer from such churn, making it difficult for analysts to actually obtain a complete list of voters who participated in a particular election [53, 44]. Nonetheless, we assume that such a complete dataset is available, for two reasons.
First, there is an important distinction between the well-known voter file, discussed above, and the lesser known, “voted file.” Maricopa County election officials retain a list they refer to as the “voted file” that is not affected by voter churn. The voted file only contains all voters that participated in a specific election and includes their contemporaneous precinct and ballot style assignment. This voted file can be obtained through an open records request. Appendix B.2 provides more detail on our comparison between this voted file and a commercial voter file.
Second, even if such a voted file is unavailable to the public, the state has access to the voted file. Any privacy concern about vote revelation should apply equally to both individual members of the public and the state itself. Some specific subtleties of the Maricopa County voted file highlight our point well. Maricopa’s Address Confidentiality program removes the records of about 4,000 voters from their public voted file. These voters are typically victims of domestic abuse and stalking who are allowed to shield their addresses from public records [[]also see https://perma.cc/AG7S-ABRH]maricopa2022. Further, Maricopa’s public voted file lumps together in-person votes and provisional votes under the same vote method. However, even though these modifications may limit the extent to which any member of the public could reveal a vote, they do not limit Maricopa County from revealing a voter’s choice, since the county retains a voter’s confidential address and full vote method. Moreover, public poll observers might be able to compile their own database of which voters cast provisional ballots to augment the voted file.
For both reasons, in what follows we measure the amount of potential revelation that would be possible with a complete list of voters and their quasi-identifiers.
Results
We first calculate the magnitude of vote revelation in Maricopa County across the continuum of election reporting and then consider the distribution of revelation by contest and by different types of voters. Finally, we compare revelation in Maricopa County to revelation in other large counties and in other states.
Prevalence of Vote Revelation
Table 2 presents our main finding, calculating whether a voter had at least one of their votes revealed in a given election. In other words, if we define $r_{ij}=1$ when voter $i$’s vote for contest $j\in\{1,...,K_{i}\}$ is revealed, and $r_{ij}=0$ otherwise, we count the number of voters for whom $r_{ij}=1$ in at least one contest $j$111111 To formalize our quantities of interest, index voters by $i\in\{1,\ldots,N\}$. The number of voters revealed in contest $j$ is given as the sum across voters, $\sum_{i=1}^{N}r_{ij}.$ In Table 2, we compute $N-\sum_{i=1}^{N}\prod^{K_{i}}_{j=1}(1-r_{ij})$, i.e., everyone except voters for whom none of their $K_{i}$ contests are revealed. In this sense, our definition serves as the worst case scenario for the magnitude of vote revelation. We restrict our analysis to contested contests because there is no concern about voter intimidation or vote buying when voters have no choice. [^12] A contested contest is defined as a contest having at least two candidates listed on the ballot, not counting write-ins, or having at least two options on a referendum (Yes - No). All referendums are defined as contested by this measure. Over 99.5 percent of the ballots cast in the 2020 general election included at least 64 contested contests.
Table 2: Revelation by Aggregation of Election Results. The table reports the number of voters with at least one revealed vote by level of reporting unit. Parentheses translate the number into a percentage of the total votes. The table focuses on the November 2020 general election and analyzes only contested single-member contests with non-transferable vote. The average voter in this data has a ballot of about 60 contested contests.
Table 2 reports how many voters could have their vote choice revealed in four types of reporting regimes, from the aggregate precinct-level to the individual ballot-level. The top panel focuses on public vote revelation while the bottom panel considers whether vote choices might instead be vulnerable to local revelation. Recall that public vote revelation is only possible where an election official reports that all voters in a single reporting unit have voted for the same candidate. In contrast, local vote revelation is possible when a reporting unit is less than unanimous. In that case, individuals may augment the reported election results with knowledge of their own vote choice such that the remaining vote is unanimous to them.
The first cell in the top panel shows that 19 voters could have at least one of their votes publicly revealed if the 2020 general election results were reported at the precinct level. This is 0.0009% of the more than 2 million voters who participated in Maricopa County’s election, or less than one-one hundredth of one-tenth of one percent. If the same quantity were reported at the precinct-method level, as in Maricopa County’s official canvass, 1,088 voters (0.05%) could have at least one of their vote choices revealed. In the most granular reporting unit we consider, where election officials release individual ballots, 3,492 voters, or 0.17% of Maricopa voters, could have at least one vote choice revealed.
Overall, the extent of vote revelation is much smaller compared to the revelation rates found in other domains. One reason why the revelation rate is small here is because there is limited linkable information in election results compared to the nominally private data in other domains. For example, [3] find that an address, date of birth, and gender uniquely identify 98 percent of Texas voters. Similarly, [29] estimates that a date of birth, gender, and zipcode uniquely identify 61 percent of the total U.S. population. But cast vote records include none of these details.
The remaining rows of Table 2 report the increase in local vote revelation relative to public vote revelation in the top row. The table uses our proposed notation of $\alpha$ to define the degree of local revelation. When $\alpha=1$, we consider the scenario in which every person knows how they themselves voted, but does not share how they voted with anyone else. In that case, we find that 56 additional voters would have at least one of their vote choices locally revealed with precinct-level aggregation. The case of $\alpha=2$ might correspond to a household within a precinct, where a couple shares their vote choice with each other. We find that 81 voters (which includes the 56 voters under $\alpha=1$) will have their vote locally revealed in this case.
Local revelation in any contest rises close to 1 percent of total votes when we consider the case of up to $\alpha=10$ individuals who share who they voted for. But the $\alpha=10$ case is implausible in practice. It requires a group of ten voters — say, 10 Republican voters in a precinct where 90 other voters voted Democratic — to find each other, credibly share their vote choices, and thereby credibly learn how the 90 other voters voted. Instead, $\alpha=10$ is best understood as an upper bound of local revelation.
Table 2 distinguishes local revelations from public revelations because the two quantities are qualitatively different. Local revelations have less serious implications for the secret ballot than public revelations. If voter $i$’s vote choice is exposed to public revelation, their vote is exposed to anyone with a voter file and election results. In other words, unanimous results reveal everyone’s vote choice to everyone. In contrast, if voter $i$’s vote choice is exposed to local revelation with $\alpha=1$, their vote choice is exposed only to that one voter in the same reporting unit. As a result, the ramifications of local vote revelation are significantly less damaging than public vote revelation.
In Appendix Table C2, we conduct the same exercise for the 2022 primary election. This primary elections was 2 to 5 times more likely to produce revelation than the general election. Given our statistical model, this is not surprising. This primary election had an electorate three times smaller than the 2020 generally election but with the same number of precincts, which mechanically increases the likelihood of revelation. Primaries may also be more likely to feature lopsided contests, which produce unanimous reporting units.
Patterns of Vote Revelation
The number of voters with at least one vote choice revealed is only one relevant dimension of vote revelation. In this sub-section, we address concerns that the harm of revelation could disproportionately affect voters in certain types of contests, who cast certain types of ballots, with certain partisan preferences, or with unpopular preferences. These patterns may be relevant to public officials and courts alike who are considering other dimensions of the secret ballot, beyond the extent of revelation.
In the analyses that follow, we focus on the distribution of public revelation in the November 2020 general election. We focus on this particular quantity, rather than local revelation or the primary election, for several reasons. First, vote patterns in Maricopa’s general election, which featured the Presidential election, are likely more generalizable to other jurisdictions. Further, Maricopa’s general election is particularly useful because of its long ballot, which allows us to compare revelation in the Presidential election with revelation in congressional, state, local, and referendum elections. Finally, as explained above, public revelation is more damaging than local revelation.
By Contest
We first disaggregate the results in Table 2 by the type of contest. Figure 5 investigates the distribution of revelation for national, state, local, and referendum contests. It then plots the percentage of potential votes revealed if results are reported at the ballot-level, with thick bars indicating the median percent revelation.131313 We drop 3 contests from the figure because they are contests with a small number of eligible voters to begin with, making the fractions across contests less comparable. For the general election, these contests are for the Gila Bend school district (613 voters, 4 potentially revealed votes), the Palo Verde education district (1145 voters, 12 potentially revealed votes), the Wilson education district (553 voters, 1 potentially revealed vote). By focusing on the ballot level, we make clear the worst-case scenario for revelation.
Figure 5: Contests with More Revelation. A boxplot showing the fraction of public revelations in the ballot-level (precinct $\times$ style $\times$ method) reporting regime. Each point represents a contest. The solid bars indicate the median, the box indicates the first and third quartile, and the whiskers extend to 1.5 multiplied by the interquartile range. Contests that are more partisan tend to have more revelation.
Overall, Figure 5 shows that top-of-the-ticket partisan offices have a higher prevalence of vote revelation than down-ballot offices. For example, the median judicial retention election has a potential revelation rate of 0.000012% in the 2020 general election, which is three orders of magnitude smaller than the comparable rate for President. As a result, concerns about the harm of revelations should generally track the salience of the contest.
These results are consistent with our statistical model. Revelations track salience because voters tend to abstain in low-salience contests like ballot measures and retention elections. For example, Appendix Figure C5 shows that the undervoting rate for the President is only 0.4 percent, while the undervoting rate for a down-ballot race such as a judicial retention election is close to 40 percent.
By Vote Method
Table 3: Revelation by Vote Method. Number of voters with at least one contest subject to public revelation in the ballot-level reporting regime, for each vote method.
We next consider revelation by vote method. Maricopa County voters can vote in person, by mail, or provisionally. Federal law generally requires that election officials offer voters a provisional ballot as a fail-safe when they cannot otherwise cast a valid ballot. Provisional ballots may be offered, for example, when an individual may not appear in a precinct’s pollbook or may not present required identification.
Table 3 shows the degree of vote revelation by vote method. As in our analysis of contests, we track public revelation if results are reported at the ballot-level. The first pair of columns shows the composition of the vote method in the sample. Conducted during the pandemic, the 2020 general election was done mostly by mail. The second pair of columns shows the composition of the vote method only among the votes revealed. Because we examine revelation at the ballot level, we can classify each ballot with revelation by the ballot’s vote method. As we saw in Table 2, 3,492 ballots have a revelation in at least one contest.
The table shows a remarkable discrepancy. Provisional ballots are only 0.3% of all ballots cast, but they are nearly half of the ballots with at least one-contest with a revealed vote. Consequently, the risk of at least one vote on a provisional ballot being revealed is 26.7%, but the risk of at least one vote on a mail-in ballot being revealed is less than 0.1%. Here, the size of the reporting unit can explain the lion’s share of the pattern across these three groups. From Table 1, we know that there are typically only 3 or 4 provisional ballots in a given precinct $\times$ ballot style combination, while there are typically over 250 such mail-in ballots.
By Voter Partisanship
A related concern might be that revelation could fall disproportionately on people with certain partisan preferences. One benefit of ballot-level data is that we can proxy for partisan preferences using the vote choice for President on each ballot.
Table 4 shows vote revelation by presidential vote choice. It counts both the number of voters who supported each presidential candidate as well as the number of such voters who have any of their votes revealed when results are reported at the individual ballot-level. Based on the two quantities, the table reports the conditional probability that a voter has any vote revealed given their presidential preference. Biden voters are most vulnerable to any-contest revelation (0.19%), while Jorgensen voters are least likely (0.10%). However, the discrepancy between candidates is nowhere near the stark discrepancy found between vote methods (Table 3).
Table 4: Revelation by Presidential Vote Choice. Number of voters with at least one contest subject to public revelation in the ballot-level reporting regime, for each presidential candidate’s supporters.
Again, our statistical model helps make sense of the differences between Democratic, Republican, and Libertarian voters. Voters for a certain candidate can be clustered by local geography or by vote method (or not at all). Votes for the third-party candidate for President are not as clustered by geography as votes for the major-party candidates: There are no pockets of neighborhoods that heavily favor Libertarians, whereas some urban precincts may heavily vote for Democratic candidates. Strong clustering implies that the ex ante probability of voting for a particular candidate in a particular unit in our model ($w$ in equation (1)), is close to the extremes, leading to a larger expected revelation. The map in Figure C6 shows that both densely populated Democratic precincts in Phoenix city center and sparsely populated Republican precincts in the outskirts of the city exhibit revelations. This explains the relative parity of Biden and Trump voters’ revelations in Table 2. Also, Figure C7 shows that this difference between Biden, Trump, and Jorgensen dissipates in ballot measures and local offices.
Revelation of Unpopular Vote Choices
Another potential concern is that the harm of vote revelation might fall disproportionately on people who hold unpopular views, regardless of their partisan preferences. Individuals whose votes are publicly revealed necessarily agree with the other voters in their reporting unit. This is because revelation only takes place when voting is unanimous within a reporting unit. However, individuals in one reporting unit may disagree with individuals in nearby reporting units. To address this concern, we compare individual voters’ revealed choices to the aggregate results in different geographic areas.
More specifically, we compare revealed voters to all other voters living in a concentric geographic circle surrounding the revealed voters’ precinct. For each fixed radius from the revealed voter, we compute the degree of agreement between revealed voters and all other voters. For example, at $x$ = 10 miles, we capture all precincts whose centroid is within 10 miles of the revealed voter (See Appendix B.3 for details). We then compute the percentage of voters in the given radius who agree with the revealed choice, and show the percentage on the vertical axis. We do this at three levels of aggregation – precinct, precinct $\times$ method, and ballot.
Figure 6: Revelation of unpopular vote choices. With each revealed vote choice, we display what percent of the voter’s neighbors share that vote choice, where neighbors are defined by a geographical distance. Solid lines show averages by candidate.
Figure 6 summarizes our agreement measure when results are reported at different levels of aggregation. The different facets focus on voters whose choice is revealed when results are reported by precinct (left), by precinct $\times$ method (middle), and by ballot (right). Each line in each facet traces the agreement between revealed voters in a unanimous reporting unit and all other voters at varying distances. Ultimately, the figure suggests that vote revelation does not disproportionately fall on people with unpopular views.
The left facet has only a single line because Maricopa County had only a single precinct whose votes for president were unanimous. In this case, all of the voters in that precinct supported Donald Trump for president. By construction, the agreement value is 100% at $x=0$ miles. At a $x=10$ mile radius, the agreement value is about 75%, meaning that 3 out of 4 voters in a 10 mile radius from the revealed voters agree with the revealed voters’ choice, while about 1 in 4 voted for Joe Biden. At a $x=100$ mile radius, reaching far outside Arizona to New Mexico, about half of individuals agree with the revealed voters.
The middle facet features more unanimous reporting units. The agreement value does not begin at 100%, even when $x=0$ miles, because one vote method within a precinct may be unanimous while another method in the same precinct is often not. [^14] In other words, because vote method is not geocoded, we consider all voters in the same precinct as this voter to be their $x=0$ radius neighbors. In fact, there are many precinct-method reporting units where the revealed vote, whether for Biden or Trump, is not supported by the majority of voters in the same precinct. That is, more than 50% of voters in a given radius disagree with the revealed voter’s choice. Yet in no case do an overwhelming number of voters disagree: the minimum agreement we see is around 25%. The solid lines show the weighted mean of all the candidate’s revealed votes. Voters for Democratic candidates tend to be more in agreement with their neighbors than Republican candidates, since Democratic strongholds are clustered in dense cities. But the solid lines all hover around 50%, indicating that the revealed vote tends to reflect the majority, not the minority, view.
The right facet displays a similar pattern. One difference between the middle facet and the right facet is the addition of revealed votes for the Libertarian Presidential candidate. Consistent with Table 4, Libertarian preferences are not clustered by geography or method. As a result, although very few Libertarian votes are revealed, those that are revealed are distinctly unpopular at any radius.
Revelation beyond Maricopa
We next expand our analysis beyond Maricopa County. Ideally, we would compare vote revelation in every state at each level of reporting, from the precinct-level to the ballot-level. However, most states currently do not report election results at the ballot-level. Instead, the precinct-level remains the most common reporting unit [6]. We thus first compare revelation nationwide at the precinct-level and then compare revelation three other large counties that report more granular election results.
Nationwide Revelation
Computing vote revelation requires knowing the number of residual votes (i.e., undervotes or overvotes) in each reporting unit. Results without residual votes are problematic because a reporting unit that appears unanimous based on votes for candidates may not be actually unanimous once residual votes are accounted for. However, the reporting of residual votes varies state by state. [6] extensively standardized and validated precinct-level data from all states. In our analysis of their data, only 488 out of 2961 counties and parishes reported election results that include the number of residual votes cast by contest by precinct.
To account for the missing data problem, Figure 7 reports the percentage of precincts that appear unanimous based on the vote for the three presidential candidates (Biden, Trump, and Jorgensen). We call these three-party unanimous precincts. The percent of three-party-unanimous precincts are upper-bounds for the extent of potential vote revelation, because they ignore potential residual votes. Any residual vote in a precinct would result in non-unanimity and thus prevent public vote revelation.
Figure 7: Proportion of Voters in Unanimous Precincts. Each number is the proportion of voters in precincts in which only one of the three 2020 Presidential candidates (either Biden, Trump, or Jorgensen) have earned votes. This is not equivalent to precincts with vote revelation, because these precincts may also have write-in votes, undervotes, and overvotes. New Mexico and Nevada are excluded because they mask the election results in small precincts, and Indiana is excluded because some of its counties do not report precinct results.
We leverage the fact that some counties in the data collected by [6] do report residual votes by precinct. We first fit a regression model predicting actual unanimity among the three-party unanimous precincts in those counties. We then apply this fitted model to three-party unanimous precincts in counties where residual votes are not reported, to assign a probabilistic estimate for whether a three-party unanimous precinct is actually unanimous (i.e., has zero residual votes). Appendix Section B.4 further explains the methodology. Through this approach, we estimate that nationwide precinct-level results could reveal 11,937 votes, if election officials included residual votes. This nationwide figure is about 20 times larger than the potential vote revelation in Maricopa County at the same level of aggregation (7 out of 2 million votes; or 0.0003%). However, total revelation is still at less than one-hundredth of one percent (0.0076%).
Precinct-method revelation in three additional counties
We also consider vote revelation in three counties — Los Angeles, CA; San Mateo, CA (Redwood City); and Cook County, IL (excluding the City of Chicago) — where, like Maricopa County, votes are reported at the precinct-method level and include undervotes and overvotes. The fraction of votes revealed in these counties is similar in magnitude to Maricopa. The fraction of total votes revealed for the 2020 Presidential race is 0.009% out of 4.3 million in Los Angeles, 0.007% out of 0.76 million in San Mateo, and only 1 voter out of 1.2 million (or 0.00008%) in suburban Cook county, compared to 0.02% in Maricopa County.
The Future of Election Reporting
Given our theoretical model and empirical results for vote revelation, we reconsider how election officials might best approach election reporting. In general, almost all election officials report election results by precinct, but they are split on whether to release individual ballots too. A typical view is that individual-level ballot records should not be released if the data could reveal at least one voter’s choice. [^15] Indeed, the North Carolina State Board of Elections will not release any cast vote records because of their interpretation of a state law which makes it a misdemeanor for anyone “who has access to an official voted ballot or record and knowingly discloses … how an individual has voted that ballot.” (N.C. Gen. Stat. §163-165.1(e)). But the state’s interpretation seems to prove too much: Table 2 shows that precinct-level election returns may also reveal a non-zero amount of vote choices, yet North Carolina still reports precinct-level results, as does almost every other U.S. jurisdiction. The distinction between vote choice revelation and voter revelation is also important. We read the North Carolina statute as prohibiting the revelation of an entire ballot, whereas the revelations we document in this study are only for particular contests. We have shown that revealing an entire ballot is much more improbable especially when the ballot has multiple contests, as in Maricopa. Based on Table 2, we consider the extent of vote revelation from ballot records to be modest relative to the number of ballots cast. But more importantly, the focus on ballot records misunderstands how all forms of election reporting can lead to vote revelation. In fact, Table 2 documents vote revelation across the entire spectrum of election reporting, from the aggregate level to the individual level. The narrow focus of the current debate thus avoids the hard question about how to balance the importance of voter privacy with similarly important value of public transparency.
In this final section, we address the future of election reporting. We first review the benefits of transparency, surveying the literature in political science. We then evaluate how some states have implemented the well-known menu of policies in the privacy literature to protect voter privacy. Ultimately, we conclude that the typical privacy-protecting measures implemented in other contexts are not good fits for the unique context of election administration. Instead, we consider new preventative approaches to reduce vote revelation while maintaining electoral transparency.
Benefits of Transparency
A growing literature in political science finds support for theories that more transparency in election reporting can reduce voter fraud or increase voter trust. Further, election transparency has become central to voting rights enforcement.
For one, granular election results might reduce fraud because they facilitate the detection of voting irregularities. In fact, the growing body of so-called election forensics uses granular election results precisely for that reason. For example, in a classic study, [66] used multiple sources of granular election results, including cast vote records, to show that the effect of poor ballot design was large enough to swing the winner of the 2000 Presidential election [[, see also]]morseworking. More recent studies similarly rely on results reported by vote method and precinct to dispel claims of election fraud [28, 21, 5] or detect actual election fraud [33].
Consistent with election forensics, international organizations monitoring elections recommend that countries publish more granular election results to reduce fraud. Two studies support the recommendation. A field experiment in the 2010 Afghanistan parliamentary elections found that notifying polling station managers that their initial vote tallies will be photographed and made public substantially reduced the manipulation of vote counts [11]. Further, an over-time study of 125 low- and middle-income countries finds that more granular election results are associated with fewer perceived irregularities [59].
More granular election results may also improve public trust by changing perceptions. [59] find that more granular results improve expert perceptions of the overall quality of international elections. Further, in the contemporary U.S., [40] shows that publishing post-election audits increases the mass public’s confidence in both the accuracy and result of the election. The effect of publicly releasing audit results is large— [40] finds that its effect on the public’s confidence in the audit is about as large as the effect of the declared winner in the audit being a co-partisan. It is possible the effect might be even larger if the ballots being audited were public, too.
Separately, granular election results have also become central to promoting minority voting rights [31, 22]. The Voting Rights Act protects racial minorities from what is known as vote dilution. In general, an electoral map may dilute the power of racial minorities only if voting is polarized along racial lines (See generally Thornburg v. Gingles, 478 U.S. 30 (1986)). The difficulty, of course, is that the secret ballot does not allow researchers to directly observe how any individual voted, never mind by race. Instead, political scientists use granular election results to estimate racial polarization and measure racial vote dilution. More granular election results allow better measurement of racial polarization and, thus, more enforcement of current voting rights protections.
Current Approaches to Reducing Revelation
Transparency, of course, has costs as well as benefits. For example, for the same reason that granular election results can promote minority voting rights by detecting racial polarization, more granular electoral results could exacerbate clientilism or facilitate retribution [60]. Broadly, there are three types of privacy policies currently used by election officials to reduce or eliminate vote revelation: some jurisdictions suppress certain election results that might be vulnerable to revelation, others aggregate small reporting units into larger reporting units, and one adds random noise to election results based on ideas related to differential privacy. So far, states have typically paired suppression with aggregation.
Suppression and Aggregation
A few states use a combination of both suppression and aggregation to address privacy concerns in election reporting. For example, Nevada, New Mexico, and Florida each report granular election results by precinct and vote method unless the size of the reporting unit is below a defined threshold. In New Mexico, “in any precinct where fewer than 5 voters voted on election day, or by absentee ballot or as early voters, the total votes in that precinct shall be reported as a single total, and not by the type of method by which voters cast their ballot” (N.M. Admin. Code 1.10.33.8). Similarly, in Nevada, “[i]f a precinct has fewer than 10” votes from either early votes or mail ballots then “[t]hose returns must be reported in combination with the regular votes of the precinct” (Nev. Admin. Code 293.357). Florida’s approach is even more specific. In Florida, “[i]f a precinct has a subtotal [for each candidate and ballot type] of one to nine votes, the Supervisor shall report zero votes in all subtotals except in the ‘Total Votes’ group for that precinct” (Fla. Admin. Code Ann. r. 1S-2.053(5)(a)(3)).161616 There are similar approaches in other countries too. For example, in Germany, any precinct with less than 50 votes must send their votes, sealed and unfolded, to another precinct where they will be mixed, counted, and reported as one. §68 ¶2, Bundeswahlordnung (Federal Election Regulation), Feb. 13th, 2020 Version. In the most recent election, less than 1 percent of polling stations were affected by this policy. https://www.buzer.de/gesetz/2263/al85770-0.htm. We thank Qixuan Yang for identifying this example.
The difficulty with each of these approaches is that they naively look at the number rather than the unanimity of votes cast. As a result, their approach is both under-inclusive and over-inclusive. A reporting unit could have more votes than the threshold and still lead to vote revelation because it is unanimous. Or it could have fewer votes than the threshold and yet not lead to revelation because it is not unanimous.
Accounting for the unanimity of a reporting unit does not resolve either the suppression or the aggregation problem. For example, suppression can never completely account for the possibility of private information. More importantly, aggregating small reporting units is not as simple as it may sound. Consider a precinct where the results for a school district are unanimous. In that case, how should election officials aggregate the results to preserve privacy? Perhaps the results could be combined with a neighboring precinct or a county-wide administrative precinct. But asking election officials to edit election reporting piecemeal can easily generate the sort of distrust that granular election results are partly meant to combat in the first place.
Random Noise (Differential Privacy)
North Carolina has charted a unique approach to add noise to aggregate election results. However, the policy draws on the same logic motivating the modern paradigm for statistical privacy, called differential privacy. That logic entails adding random noise to the results of database queries (e.g., how many votes did Joe Biden receive) in order to keep sensitive information (e.g., who did a particular voter vote for) private. Data users essentially are told the underlying variance of the random number generator but only observe one draw of it, so the sensitive attribute is kept private.
Adding random noise is a powerful privacy protection mechanism for sharing sensitive data because it is more resistant to background knowledge attacks, while preserving some ability of the social scientists to draw correct inferences about the structure of the underlying data [20, 23, 1]. For example, by releasing the noised data of true, sensitive information, the decennial Census allowed outside researchers to report the bias and variance of their privacy-protection mechanism [42]. However, a user’s ability to reliably recover this summary statistic requires transparency about the method of adding random noise. Unfortunately, the North Carolina Board of Elections has yet to publicly document its random noise algorithm. As a result, while its process may well provide increased privacy protection, it does so at the cost of making it impossible to know if reliable answers to many questions can be drawn from the electoral data that it publishes.
Despite the advantages of differential privacy over suppression, this paradigm faces important conceptual challenges when it comes to election reporting. A core challenge is that certain queries of an election database (such as which candidate received the most votes) must be reported with perfect fidelity. Classic differential privacy from [19] requires that the reported result of every query to a confidential database be effectively equally likely regardless of whether any individual is removed from the database. The most important query for election results is simple: which candidate won the most votes? Differential privacy, though, requires that the answer to this query not substantially depend on any one vote, or, put another way, that we pick election winners in a way that no single vote could ever matter. That is why only stochastic manipulations, i.e., those that add some sort of random noise, can be differentially private with a finite private budget.171717 To formalize, consider the statistic $s$ (e.g. the winner of the election) produced from a private dataset of voters $D$. The (usually random noise-inducing) processing is denoted $M$, and the value of the statistic is $m$. The privacy literature considers $M$ to be differentially private if the ratio of $\Pr(M(s,D)=m)$ to $\Pr(M(s,D^{\prime})=m)$ is below a user-specified constant $\exp(\epsilon)$, where $D^{\prime}$ is a dataset that differs from $D$ by the inclusion of one additional row. Now consider an election is decided by one vote, $D$ is the data set that includes every voter and $D^{\prime}$ is the data where one voter is excluded. Then if $M$ has no noise, $M(s,D^{\prime})$ will report the same winner as $M(s,D)$ with probability 0. Therefore, the ratio of $\Pr(M(s,D)=m)$ to $\Pr(M(s,D^{\prime})=m)$ will be infinity, and no value of $\epsilon$ would satisfy the DP criterion. For $\epsilon$ to approach 0, the result of the election would have to be noised enough that the outcomes of the closest elections would be determined by lot and the presence or absence of any one ballot could never matter.
Another conceptual challenge is that classic differential privacy is defined on multiple hypothetical versions of the same dataset, comparing whether one person is in the dataset or not. This idea may work for a survey sample, where there may be multiple samples from the same population, or for a policy of continuously releasing statistics from a changing population [68]. However, an election only happens once. In fact, elections are defined by their finality. We argue it is therefore difficult to rely on a notion of multiple versions of the same election.
Many of these challenges might be solved by advances in differentially private algorithms. Some recent variants of differential privacy do incorporate these constraints into the algorithm as starting constraints so that, in our case, the certified election totals at the constituency level would not be altered [27, 16]. But these methods are complex, require expertise and documentation, and are yet to be implemented in election administration.
Preventive Approaches
The methods to eliminate vote revelation after an election, surveyed above, all have their shortcomings. Some may not provide protection from sophisticated privacy attacks [14]. Most demand the time and resources of election officials when both are scarce. And none can protect the secrecy of the ballot in terms of vote revelation from the election officials themselves.
In contrast, several preventive approaches would at least reduce or perhaps obviate the need for post-hoc approaches. These approaches will take more time to implement, but may be more feasible.
First, redistricting could reduce vote revelation. Drawing district lines for multiple offices to avoid rare ballot styles would increase the size of reporting units and thus reduce the extent of revelation. Figure A1 in the Appendix makes clear that reporting units that contain about 50 voters would drastically reduce the number of potential revelations. Yet a considerable amount of reporting units — at least one in ten ballot style $\times$ method combinations in Maricopa — are currently cast by 5 or voters or fewer (Table 1). Of course, some ballot styles and precincts may be small for valid reasons, such as making polling stations available in low-density areas or maintaining equipopulous districts. But other motivations that lead to rare ballot styles, such as gerrymandering for incumbent protection, may not be worth the costs to privacy.
Another preventive approach is to modify the information available in the voter file. As the Venn diagram in Figure 1 shows, voters’ names and addresses are easily linked to unanimous election results because names and addresses are publicly available in U.S. voter files, along with information on whether registrants turned out to vote. Limiting such information on the voter file would make vote revelation much more difficult. But it would also threaten to reduce political parties’ ability to turn out the vote [35]. A similar but less drastic approach would limit the snapshots of the voter file made available, in effect using the churn of the electorate for privacy protection.
A related preventive approach could focus specifically on provisional ballots. More than half of all revelations from releasing individual ballots are from provisional ballots (Table 3). Policies that reduce the use of provisional ballots would thus go a long way towards protecting privacy. Further, while there is a clear benefit to identifying voters whose provisional votes need to be cured, once those ballots are counted, the benefit to identifying which ballots were provisional ballots is less clear. Cast vote records could thus instead identify whether a provisional ballot was cast by mail or in person, rather than provisionally. In fact, that is exactly what Maricopa does in their voter file. The upshot would be larger reporting units and thus less revelation.
Conclusion
The upcoming presidential election will likely highlight the difficult tradeoff between transparency and privacy in election reporting. Armed with the technological ability to release individual, anonymous ballots, some election officials will now do so. But this action might reveal individual vote choices. To inform the ongoing debate, we presented the first comprehensive analysis of how the release of election results, from the aggregate precinct-level to the individual ballot-level, could lead to what we term vote revelation.
Using Maricopa, Arizona, as a case study, we find that cast vote records could reveal the vote choices of only about two-tenths of one-one hundredth of voters. While our findings cannot provide an absolute threshold to judge the extent of vote revelation, we interpret such revelation as small enough to make the benefits of transparency worth the costs of disclosure. Further, the extent and distribution of vote revelation may not determine whether a practice violates any particular state’s right to a secret ballot. That is because the legal right may be broader than our quantitative definition. Nonetheless, we believe our theoretical model and empirical evaluation inform the emerging debate about how to best address the secret ballot in a time of voter distrust.
The importance of transparency in elections should serve as a serious counterweight against suppressing, coarsening, or adding random noise to reported election results. But the importance of transparency also means that the release of election results should be robust to bad faith actors. Across the country, election officials have been the target of a coordinated campaign by election denialists to request cast vote records and other detailed election data [[, see, e.g.]]surryPropublica. While some records requests are made in good faith, the barrage of records requests has hindered the work of local election officials, who are often already at capacity [30]. To take transparency seriously, we must support election officials’ capacity to maintain and release fine-grained election results.
Appendix
Appendix A Statistical Model of Vote Revelation
We are interested in the number of publicly revealed votes in a reporting unit, which we denote by the random variable $R.$ The reporting unit contains $N$ voters and we consider a single contest with $H$ candidates. In the main text, we showed that when $H=2$ and with no abstentions, the expected number of revealed votes $\mathrm{E}[R]=N\left(w_{1}^{N}+(1-w_{1})^{N}\right)$ where $w_{1}$ is the a priori probability that a randomly selected voter in the reporting unit votes for candidate 1. In this section, we provide a more general characterization of $\mathrm{E}[R].$
Public Revelation
Let $s$ be the a priori probability that a randomly selected voter in the reporting unit abstains (undervotes or overvotes), and allow for more than 2 candidates. We again denote $w_{h}$ to be the probability that a randomly selected voter in the reporting unit will support candidate $h\in\{1,\dots,H\}$ conditional on not abstaining. Then the expected number of revelations is given by:
$\mathrm{E}[R]=N\left(s^{N}+(1-s)^{N}\sum_{h=1}^{H}w_{h}^{N}\right).$ (A1)
In the special case when $s=0$, this reduces to $\mathrm{E}[R]=N\left(\sum_{h=1}^{H}w_{h}^{N}\right).$ This quantity is small when $s$ is large and the vector of $w$ is equally distributed across $H$ candidates.
The solid line in Figure A1 shows how the expected number of public revelations ($\alpha=0$) changes as $N$ increases for different values of $\alpha$ in each of two different electoral contexts. In the left panel, we consider a lopsided contest with two candidates, $w_{1}=0.95,$ $w_{2}=0.05$, and low abstention, $s=0.05$. In the right plot, we consider a contest in which support is expected to be evenly split race across four candidates, $w_{1}=w_{2}=w_{3}=w_{4}=0.25,$ with higher abstention, $s=0.20.$ For a given $N$, the left plot with low abstention and a lopsided contest shows a larger expected number of revelations than the right plot with high abstention and a large field of candidates. Notice that while the expected number of revelations at first increases in $N$ in some scenarios, it eventually decreases in $N$ in all scenarios. Also notice that in every scenario considered in the Figure, the expected number of revelations is increasing in $\alpha$, all else equal.
Figure A1: Determinants of Revelation. Expected revelations according to our statistical model. The horizontal axis shows the number of voters. In the left panel, we show simulations from parameters that are conducive to revelation; the right panel shows the opposite. The main text includes parameter details.
Local Revelation
We next consider local revelation at different values of the parameter $\alpha>0$, where $\alpha$ denotes the number of otherwise known vote choices. We denote this quantity as $R^{\alpha}$ to distinguish it from public revelation.
In the model, the probability of revelation can be calculated from the distribution function for the maximum frequency of a multinomial distribution with parameters $\pi=(s,(1-s)w_{i},(1-s)w_{2},...,(1-s)w_{H})$ and number of trials $N$. Revelation occurs if the largest frequency is at least $N-\alpha$. For example, let $N=100$, $\alpha=1$, and $H=5.$ There is $\alpha=1$ local revelation when $100-1=99$ voters choose the same alternative (candidate or abstain), another alternative gets 1 vote, and the three other alternatives get 0 votes. Thus, we are interested in the probability that the maximum of the votes counts across the alternatives is at least 99. This is a well-studied quantity in probability theory. When $\alpha<N/2$, it is given by
$\mathrm{E}(R^{\alpha})=(N-\alpha)\Bigg{[}&\underbrace{\sum_{m=N-\alpha}^{N}{N\choose m}s^{m}(1-s)^{N-m}}_{\text{Prob. of}~{}N-\alpha~{}\text{or more abstentions}}\\ &+\underbrace{\sum_{m=N-\alpha}^{N}\sum_{h=1}^{H}{N\choose m}((1-s)w_{h})^{m}\left(1-(1-s)w_{h}\right)^{N-m}}_{\text{Prob. that some candidate received more than}~{}N-\alpha~{}\text{votes}}\Bigg{]}.\end{aligned}$ (A2)
where (${N\choose m}$ indicates the combinations for formula (binomial coefficient) $N$-choose-$m.$ In the special case when $s=0$, then the quantity reduces to
$\mathrm{E}[R^{\alpha}]=(N-\alpha)\left(\sum_{m=N-\alpha}^{N}\sum_{h=1}^{H}{N\choose m}w_{h}^{m}\left(1-w_{h}\right)^{N-m}\right).$
All else equal, revelations are smaller when $N$ is large. Revelations increase as $\alpha$ grows. Figure A1 shows how revelations differ when $\alpha=1$ and $\alpha=2$. Note that we multiply the probability of revelation occurring by $N-\alpha$ to get the expected number of revelations because $\alpha$ of the $N$ votes are not revealed by the election data, but rather by the voters themselves. The election data can only reveal the remaining $N-\alpha$ votes.
When $\alpha\geq N/2,$ there is no compact expression of the number of revelations. The reason is that as long as $\alpha<\frac{N}{2}$, it is not possible for two or more alternatives to receive a number votes that allow for revelation at the same time. When this condition holds (which it always does when $\alpha=0$) then we can find the chances of revelation by summing up the (disjoint) probabilities that each candidate would receive $N-\alpha$ or more votes. However, for larger $\alpha,$ the total probability of revelation can no longer be expressed as a simple sum of the probabilities that each alternative will receive the number of votes required for revelation. The dotted lines in Figure A1 show how local revelation is strictly more than public revelation, all else equal.
Empirical Results
Our theoretical model illuminates the main components of vote revelation. We do not, however, recommend applying our model to predict actual vote revelation. There are two difficulties in applying our model to predict the magnitude of revelation in a given reporting regime. First, the vote probabilities $s$ and $w_{h}$ are not known ex ante. It may be easy to predict the aggregate vote share for a voter, but note that these values are defined at the reporting unit level in our model, rather than statewide. Second, predicting the implications of increasing $\alpha$ in a real-world setting faces the challenge that $N$ also varies by reporting unit. As our summary statistics show, many granular reporting units have only 1 or 2 voters, while some have thousands. For these reasons, the thrust of our study focuses on computing potential revelation from actual data.
Appendix B Data Construction
Cast Vote Records
Figure B2 shows an example of an actual ballot image and its cast vote record representation. This particular ballot is a rare example in which the voter mistakenly signed their name on the ballot (redacted at the top of the document with a black rectangle). Through an exhaustive search, [4] find that 22 out of the 118,216 ballots (0.02%) cast in Leon County, Florida’s 2022 general election include a signature that may or may not indicate the voter’s name. The code that indicates the precinct is also redacted (near the bottom of the ballot) because of the state’s rules to redact precinct from units where at least one vote method has fewer than 30 votes (Fla. Stat. §98.0981(3)(a)1. (2023)).
Figure B3 illustrates the distinction between a precinct and a ballot style. The figure depicts two political jurisdictions, a city and a legislative district, which are not coterminous. While all city residents live in the first legislative district, suburban residents are split between two different legislative districts. The result illustrates the complicated geography of election administration in the U.S. In the city, the two precincts will both share the same ballot style (precincts A and B sharing ballot style X), since everyone in the city is in the same municipal and legislative district. In contrast, in the suburbs, a single precinct will be split by two ballot styles c(ballot styles Y and Z in precinct C), since some suburban residents reside in the first legislative district while others reside in the second. The result is three distinct ballot styles.
Figure B2: Examples of Ballot Images and Cast Vote Records. (a) An image of an actual ballot from [4]. This particular ballot is a rare example in which the voter mistakenly signed their name on the ballot (redacted at the top of the document with a black rectangle). (b) A mock example of a cast vote record of the same ballot. Information about the precinct, method, and vote choices are stored, but not marks such as the signature. The precinct and the ballot identifiers are hypothetical values.
(a) Ballot Image
(b) Cast Vote Record representation
Figure B3: Precincts and Ballot Styles. The figure depicts our definition of precincts and ballot styles in an example of an area split by two district lines.
Voter File and Voted File Data
Revealing votes requires not only information about the ballots cast, but also about the individuals who cast those ballots. Thus, in addition to studying what would be revealed if cast vote records or granular election returns are combined with a complete and accurate enumeration of the voters casting ballots in the election, we also investigate the availability of such an enumeration. As depicted in Figure 1A, vote revelation requires an accurate turnout record for every voter in a unanimous reporting unit.
We collect two types of individual-level turnout records, both based on administrative data on voter registration and participation: a commercial voter file produced by L2, a private data vendor, and an administrative file called the “voted file” produced by Maricopa County.
The L2 voter file identifies a voter’s name, birthdate, and address, as well as whether or not each voter turned out in a specific election and by what voting method. However, consistent with the churn in the electorate, the January 2021 L2 voter file contains 26,819 fewer voters than the votes cast in the November 2020 elections (about 1.3% of the total votes actually cast).
The voted file, technically called the VM55 in Maricopa, is a complete record of the voters in each election. It includes precinct and a coarse measure of vote method. Although a voter’s ballot style is not included in the voted file, it is possible to map each voter’s known registration address to a ballot style with publicly available precinct maps. The only exception to this is if the ballot style in question is a federal-only ballot. Using the 2020 VM55, we verified that the voted file from Maricopa county matches the number of ballots obtained in the cast vote records, with one exception of the protected voter registration discussed above.
As discussed in the main text, we put aside the limitations of the voter file data and focus on potential vote revelation, i.e., assuming that the actor seeking to reveal votes has access to an accurate enumeration of voters that indicates the ballot style, precinct, and method of voting for each voter. Accounting for inaccuracies in the enumeration of voters participating in the elections would reduce the amount of vote revelation. However, it is possible that someone seeking to reveal votes could produce a more accurate enumeration of the voters than we have access to, so we take the conservative approach of reporting the degree of revelation that would be possible if the best possible enumeration of voters was available.
Revelation Agreement by Distance
This section describes some details for constructing the data underlying Figure 6. Figure B4 shows one example of the procedure described in the main text. The blue circle indicates a 10 mile radius. Figure 6 shows that the value of at $x=10$ in the left panel is about 65 percent. That means that, among voters whose precinct’s centroids are in the blue region of Figure B4, about 65 percent of them are Trump voters. As the radius widens, values necessarily converge towards the national vote.
Figure B4 only shows the Arizona map, but we combine shapefiles from neighboring states using [64] to expand our measure across state borders.
Figure B4: Defining Voter’s Neighbors. The blue circle shows a 10-mile radius circle around the center of Maricopa’s Sky Hawk precinct. Sky Hawk precinct, shown in red, cast all 31 of its in-person votes in 2020 Presidential election for Donald J. Trump. To estimate the rate at which voters living within 10 miles of the center of Sky Hawk precinct also supported Trump, we identify those precincts with centroids that are within 10 miles of Sky Hawk’s center. Those precincts are shaded blue in the figure. We then calculate Trump support across those precincts.
National Election Data
We used the following procedure to estimate the total number of voters across the whole country whose Presidential vote would be revealed if results were reported at the precinct level. The main idea of our modeling is as follows. We focus on precincts that appear unanimous in jurisdictions that do report residual votes and consider if they are actually unanimous when accounting for residual votes. We fit a logistic regression where the outcome is 1 if the apparently unanimous precinct also had 0 write-in and residual votes (i.e., in fact unanimous), and 0 otherwise (i.e., some voter also voted write-in, undervoted, or both). Using the regression, we then generate predicted probabilities that each apparently unanimous precinct in the nationwide data is actually unanimous.
  1. (a)
[…]
  2. (b)
    Subset the data to jurisdictions that report residual votes (including those imputed above) and write-in votes. This limits the data to precincts in 487 counties.
  3. (c)
    Subset the data further to precincts in these jurisdictions that are 3-party unanimous: precincts in which only one of the three 2020 Presidential candidates – Biden, Trump, and Jorgensen – have earned votes. This limits the data to 173 precincts in 46 counties (in the following states: AL, FL, IL, NY, OR, TX, WY). 50 percent of these precincts are Trump precincts in which Trump won positive votes and Biden won none. Figure 7 shows the voters in such 3-party unanimous precincts as a fraction of the total votes in the entire state.
  4. (d)
    Fit a logistic regression in this data, where the outcome is 1 if the precinct also had 0 write-in and residual votes (i.e., in fact unanimous), and 0 otherwise (i.e., some voter also voted write-in, undervoted, or both). The predictors in the regression are the log of the total votes for the three candidates, and whether the precinct is a Trump district. About 96 percent of the precincts in the data have an outcome of 1. The logit coefficients are presented in Table B1.
  5. (e)
    We then return to the national data from (a), and subset it to 3-party unanimous precincts. We therefore take data regardless of whether it reports residual votes by precinct. This results in 1740 precincts. We then generate predicted probabilities from (d) for each precinct, generating the probability that that precinct is actually revealed.
We take the weighted sum of each precinct’s vote count in (e), weighted by the predicted probabilities. This sum generates 11,139. This amounts to 0.0071% of the total number of votes reported in the 50-state MEDSL Data.
Table B1: Model predicting unanimous results among precincts where all candidate votes were for one candidate. Table shows logit coefficients and standard errors in parentheses. Unanimous results are defined here as having voted for a single Presidential candidate with no overvotes or undervotes. The variable “support for Trump” distinguishes all-Trump precincts or all-Biden precincts.
Outcome: Actual Unanimity
Intercept 3.62
log(Total votes for Biden/Trump/Jorgensen) -0.41
Support for Trump 0.27
Mean of Outcome 0.96
N 173
AIC 62.0
Appendix C Additional Figures and Tables
Table C2 shows the prevalence of revelation in the 2022 primary election. Primary elections have a smaller electorate and smaller reporting units, which increases the risk of revelation. On the other hand, voting patterns for precinct elections may be less lopsided at geographic levels. Similar to Table 2, we show revelations for at least one contest by level of aggregation. At the precinct level, 68 primary voters have their vote revealed in at least one primary contest, compared to 19 in the general election. At the ballot level, 2,823 primary voters are likely to have their vote revealed, while the number is 3,392 in the general election. In that case, 0.33% of Maricopa primary voters could have their vote choice revealed.
Figure C5 shows summary statistics of the contests used in Table 2 and Figure 5. The vote for President was one of the most competitive races countywide with the least amount of undervotes, while nonpartisan judicial retention are lopsided but also feature high amounts of undervoting.
Figure C6 shows where revelation is most likely to occur. The left map colors each precinct by the number of ballot-level revelations, and the right map colors each precinct by the Presidential voteshare. Revelations occur in both sparsely populated rural areas (that tend to vote Republican) and densely populated urban areas (that tend to vote Democratic).
Figure C7 uses the classification of Trump, Biden, and Jorgensen voters in Table 4 and counts the number of revelations for each group of voters. It shows that Biden voters and Trump voters are more likely to have their votes revealed in the US Senate race and US House race, but that pattern dissipates in lower-level election.
Table C2: Revelation by Aggregation of Election Results. Follows Table 2 but showing the 2022 Primary Election rather than the 2020 General Election.
Figure C5: Voting Patterns by Contest. Each point is a contest in the Maricopa 2020 November election, characterized by the level of undervoting (no valid vote for any particular choice) on the horizontal axis and the level of lopsidedness on the vertical axis. The figure only uses contests with two official choices and where the voter makes only one vote.
Figure C6: Precincts with Vulnerability of Revelation. The left panel shows the number of public revelations for President if results were reported at the ballot level. Precinct boundaries come from the Maricopa’s 2020 shapefile. The right panel shows the Biden-Trump voteshare in each precinct using data from [64]
Figure C7: Contest-Specific Vote Revelation by Presidential Vote Choice. Boxplots omit all outliers, only show median + IQR + whisker.